### v1

# TIPO: Text to Image with Text Presampling for Prompt Optimization

Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation.
Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering. The previous works [31] tried to utilize LLMs to lenghthen prompt to solve tendency, however many ignore the root cause of why some prompts result in worse images and rely heavily on manually engineered prompts, which require substantial resources.
To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
Unlike previous approaches that focus on extending prompts with manually collected prompt sets [4, 15, 45, 21], directly use LLM to modify the prompts [26, 51] or utilize reinforcement learning like method which update model’s input or model itself from text-to-image feedback [22, 13, 31] without addressing underlying issues, TIPO introduces a method that optimizes prompts efficiently as shown Section 4.
Our work makes the following key contributions:
  1. 1.
    We introduce an universal prompt optimization framework that is designed to work with any text-to-image model, providing a versatile solution for improving image generation across different architectures.
  2. 2.
    We provide comprehensive evaluation methodology and results including using both standard and non-standard metrics. Our experiments demonstrate that TIPO’s prompt engineering method can improve image quality in terms of mathematical measures and user preferences, validating its effectiveness from multiple perspectives.
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompt refinement but lacks understanding of T2I-specific prompts. (b) improves detail but limits variety. (c) optimizes for model feedback but are computationally expensive. (d) TIPO (ours) aligns prompts with the trained distribution of T2I models, ensuring prompt detail, consistency with training data, and controlled variety within model scope.
This paper presents TIPO’s theoretical foundations, implementation details, and experimental results demonstrating its effectiveness in improving aesthetic scores [17], reducing image corruption [32, 33], and better aligning generated images with dataset distributions [24, 35]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
2 Related Works
2.1 Direct Prompt Refinement Through LLMs
With the advent of advanced LLMs, several projects have leveraged their few-shot in-context learning capabilities to optimize prompts for Text-to-Image (T2I) models. For instance, CogView3 [51] and the work by Lee et al. [26] utilize GPT-J and Text Style Transfer (TST) techniques for prompt refinement. However, this approach relies heavily on the LLM’s inherent knowledge of visual content descriptions, which can be unpredictable and may not always align with specific T2I model requirements.
2.2 Extending Prompts with Manually Collected Data
In the open-source T2I community, a common approach involves using curated datasets of “good prompts” to fine-tune or train LLMs for prompt extension [4, 45, 15]. These models are typically trained on image-text pairs collected from users of proprietary T2I services like Midjourney [^1] https://www.midjourney.com or Dall-E 3 [6]. While this method is straightforward and somewhat effective, it often lacks the versatility to adapt to diverse T2I models beyond the one used in the original dataset collection.
2.3 Image Output-Based Tuning
Recent works have explored more sophisticated prompt optimization techniques based on image outputs.
Reinforcement Learning Approaches: Promptist [22] introduced a two-stage framework combining supervised fine-tuning with reinforcement learning (RL). The RL stage incorporates aesthetic and relevance scores as environmental feedback, enabling the model to learn optimal prompt refinement strategies.
Iterative Optimization: OPT2I [31] proposed a backpropagation-free prompt optimization framework. This method uses scorers and LLMs to iteratively update input prompts, aiming to improve aesthetic quality and coherence.
These RL-based or RL-like approaches allow for prompt refinement tailored to specific model properties or user preferences. However, they often require extensive computational resources and may not generalize well across different T2I models or datasets.
While these methods have shown promise, they often face challenges such as difficulty in maintaining alignment with original user intent, challenges in collecting manually engineered prompts, high computational costs for RL-based methods, and limited generalization across diverse T2I models. By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
Figure 2: Comparison of prompt optimization methods for a T2I model [7]. (a) yields a basic image with limited detail. (b) provides specific output but requires manual engineering. (c) introduce nonsensical elements (red boxes) in the teal region, where the model’s actual output distribution for a given prompt (green) and all the image distribution (blue) overlap outside the user’s expected output distribution (red). (d) captures user intent accurately, balancing detail and diversity.
3 Prompt Optimization
3.1 Motivation and Challenges
Nowadays text-conditioned models [7, 37, 19] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive. Moreover, simply creating a longer prompt does not guarantee improved results; it must be crafted in line with the model’s specific prompting conventions and interpretive nuances.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the model’s familiar vocabulary and stylistic tendencies [6], while others rely on direct feedback from the model to iteratively improve prompt alignment [22]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text inputs. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
3.2 Formulation
To formalize prompt extension, we define $\mathcal{P}$ as the set of all possible prompts, $\mathcal{N}$ as the Gaussian noise vector space, and $\mathcal{I}$ as the set of all possible images. A T2I model then acts as a mapping function:
$f(p):\mathcal{N}\rightarrow\mathcal{I}_{p}$
where, for any prompt $p\in\mathcal{P}$, the model maps noise vectors from $\mathcal{N}$ to a subset of images $\mathcal{I}_{p}\subseteq\mathcal{I}$ associated with the prompt $p$.
Now we define a pair of prompts: simple prompt $p_{s}$ and detailed prompt $p_{d}$. Formally, for a simple prompt $p_{s}$ and a detailed prompt $p_{d}$ that extends $p_{s}$, we have the following relationship:
$f(p_{d})(\mathcal{N})\subseteq f(p_{s})(\mathcal{N})$
This formulation implies that the image subset generated by the detailed prompt $p_{d}$ is contained within the subset generated by the simple prompt $p_{s}$, thereby preserving the relevance of the initial prompt while adding specificity.
To further encapsulate the prompt generation process, we define the prompt extension function, $\mathcal{E}$. This function takes a simple prompt $p_{s}$ and additional metadata $\mathcal{M}$ (such as tags) as inputs, and produces a detailed prompt $p_{d}$ as output:
$\mathcal{E}:\mathcal{P}_{s}\times\mathcal{M}\rightarrow\mathcal{P}_{d}$
where $\mathcal{P}_{s}$ is the space of simple prompts and $\mathcal{P}_{d}$ is the space of detailed prompts. Here, $\mathcal{M}$ provides auxiliary information that guides the extension of $p_{s}$ into a more refined prompt $p_{d}$.
For each instance of prompt extension, we apply the function as follows:
$p_{d}=\mathcal{E}(p_{s},M)$
where $M$ could represent a specific metadata component, such as $T_{s}$, or other forms of contextual information depending on the input type.
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory (and are not simply creative combinations). Thus, while random word extension typically yields unreliable results, it may occasionally produce a usable prompt by chance. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\supseteq\mathcal{I}\supseteq\mathcal{I}_{p_{s}}$
Which means:
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
This scenario is illustrated in Figure 2, where “Simple Prompt + Random Words” shows a scattered distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework
4.1 Overview
TIPO introduces a dataset-driven approach to prompt optimization by training directly on the existing dataset. This avoids reliance on manually engineered prompts [4, 45, 15] or on computationally intensive RL techniques [22, 31]. Methods using LLMs rely on general language knowledge, which may misalign with specific T2I model needs [6, 51]. In contrast to previous approaches, TIPO captures the model’s training data distribution closely.
4.2 Distribution-Aligned Prompt Extension
Central to TIPO’s method is an extension function $\mathcal{E}$, which iteratively refines prompts to enhance both the diversity and quality of generated images while remaining aligned with the intent of the original user-provided prompt and the training data distribution. By applying $\mathcal{E}$ to a simple prompt $p_{s}$ to generate a detailed prompt $p_{d}$, TIPO aims to approximate the target distribution:
$\bigcup_{i=0}^{n}f(p_{t_{i}})(\mathcal{N})\approx\mathcal{I}_{p_{s}}$
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
4.3 Prompt Structure
The construction of simple prompts, $p_{s}$, and their corresponding detailed prompts, $p_{d}$, varies depending on the dataset and input modality. We consider two primary scenarios: tag-based captions and natural language (NL) captions.
Tags are used to describe the content of images and are commonly formatted as a comma-separated string when input into T2I models. For a given image-tag pair, let $T_{n}=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ represent the complete set of tags for an image. We derive the simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ where $m<n$, and define the detailed tag set as $T_{d}=T_{n}$. The simple and detailed tag prompts are then constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{d}=\text{concat}(T_{d})$
NL captions are typically generated by LMs with vision modality [29, 1, 2, 27, 5, 14, 48, 16] and may contain multiple sentences. Each image is associated with either two types of NL captions—a short NL caption $S_{s}$ providing a brief, general description, and a long NL caption $S_{d}$ offering an extended, detailed description—or a single long caption that can be segmented. For images with a single long caption, let
$S_{d}=\{\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}\}$
represent the complete set of sentences in $p_{d}$. We derive the simple sentence set
$S_{s}=\{\text{sentence}_{1},\text{sentence}_{2},\ldots,\text{sentence}_{m}\}$
where $m<n$, and define the detailed sentence set as $S_{d}=S_{n}$. The simple and detailed NL prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{d}=\text{concat}(S_{d})$
These ensure that $p_{d}$ is a superset of $p_{s}$, aligning with the concept that detailed prompts provide a more specific description.
To generate $p_{d}$ from $p_{s}$, TIPO employs two distinct formats based on the relationship between $p_{s}$ and $p_{d}$:
  1. 1.
    When $p_{s}$ is a substring of $p_{d}$: The format <meta> <p_d> is used. As TIPO model is an autoregressive LM, $p_{d}$ is directly used after the <meta> tokens, allowing the model to generate subsequent tokens based on any substring of $p_{d}$.
  2. 2.
    When $p_{s}$ is not a substring of $p_{d}$: The format <meta> <p_s> <p_d> is employed. Since $p_{s}$ is not contained within $p_{d}$, both are included separately after the <meta> tokens to guide the generation process effectively.
4.4 Prompt Formatting
In TIPO, we define a simple and consistent format for the content of the <meta> token or the simple prompt $p_{s}$. The format is as follows: <Category>: <Content>. Metadata categories might include artist, copyright, aspect ratio, quality, and year. For instance,
quality: masterpiece, year: newest
Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
4.5 An Example Scenario and Workflow
Figure 3: (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from user input (short tag prompt $p_{s}$) to final output (detail prompt $p_{d}$). The shading from gray to light sky blue represents an increase in context richness in the prompt.
Figure 3 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence. During each generation cycle, one input type is treated as the primary prompt which is target prompt for extending or refining, while the other types are used as <Content> of metadata to enrich the context. Figure 3 provides an example of this setup, where:
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\\ \end{array}$
The generation proceeds sequentially as follows:
  1. 1.
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$
  2. 2.
    tag_to_long: $T_{d}$ is incorporated into the metadata, and TIPO produces a refined short NL prompt $S_{s}$ based on $T_{d}$.
  3. 3.
    short_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
  4. 4.
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
This example highlights how TIPO leverages an autoregressive LM to perform prompt extension in a stepwise, sequential processes, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle. More detailed task definitions for TIPO can be found in Section 5.2. By iteratively refining prompts in this structured way, TIPO addresses the limitations of simplistic or randomly extended prompts, enabling it to generate high-quality, diverse images that closely align with user intent.
4.6 Versatile and Scalable Design of TIPO
The example in Section 4.5 demonstrates how TIPO’s structured approach enables it to create coherent, intent-aligned prompts. Building on this foundation, TIPO’s design extends smoothly to a wide range of models and datasets, showcasing its versatility and scalability.
A single TIPO model trained on dataset D can be applied to all text-to-image models trained on the same dataset, enabling broad applicability across different architectures and versions that share the same training data. Similarly, a TIPO model trained using caption method M (e.g. tag, NL captions) can be used with any T2I model trained with captions from method M. Experimental results in Section 6 further illustrate this adaptability, as TIPO successfully applies across multiple model architectures and training configurations, confirming its broad applicability and robustness.
5 Training TIPO model
In this section, we detail how to build the TIPO model based on the TIPO framework. For various T2I models [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7], tag and NL captions are widely used captioning methods. To allow TIPO model to seamlessly transform captions between different formats or refine captions within the same format, we define specific tasks that enable flexible and precise prompt construction.
5.1 Modeling
TIPO adopts a causal autoregressive LM framework as prompt extension function, $\mathcal{E}$, training to extend and refine user-provided prompts into detailed, content-rich prompts $p_{d}$ from simpler inputs $p_{s}$, thereby effectively balancing specificity with alignment to user intent. This process captures a subset of outputs closely aligned with the user’s original intent, ensuring relevance. Through use of existing datasets for prompt optimization, along with its iterative extension function and application of LM, TIPO achieves significant coverage and diversity in generating high-quality images aligned with training distributions, all without the need for manual prompt engineering.
5.2 Task Definitions
TIPO encompasses three primary tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. To facilitate these tasks, we define several specific task types and corresponding formatting methods during training:
  1. 1.
    tag_to_long: Use tags as metadata to generate a new NL prompt or extend a user-provided NL prompt.
  2. 2.
    long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  3. 3.
    short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
  4. 4.
    short_to_long: Use a user-input NL prompt as metadata to generate a refined detailed prompt $p_{d}$.
  5. 5.
    short_to_tag_to_long: Use a user-input NL prompt or tag sequence as metadata to generate a refined detailed prompt $p_{d}$.
  6. 6.
    short_to_long_to_tag: Use a user-input NL prompt or generated NL prompt as metadata to extend a tag sequence.
  7. 7.
    tag_to_short_to_long: Use user-input tags or NL prompts as metadata to generate a refined detailed prompt $p_{d}$.
By defining these task types, we can achieve special cases where, for instance, inputting a short description or a short tag sequence results in a full-size tag sequence and NL prompt in a single pass (i.e., generating $T_{d}$ and $p_{d}$ simultaneously) as shown in Section 4.5.
5.3 Model and Training Details
We utilize the LLaMA architecture [46, 47] with models of 200 million and 500 million parameters:
  • TIPO-200M: Pretrained on the Danbooru2023 [34, 49] and GBC10M [25] datasets for 5 epochs, then fine-tuned on the Danbooru2023, GBC10M, and CoyoHD11M [9] datasets for 3 epochs, resulting in a total of approximately 40 billion non-padding tokens seen.
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, resulting in a total of approximately 30 billion non-padding tokens seen.
Note that we calculate the token count based on non-padding tokens only. Given the relatively short and variable lengths of data entries, the effective token count may be lower than anticipated.
5.4 Training Procedure
In our experiments, each training pass generates a new prompt of a single type. To cover various prompt types—such as a refined prompt $p_{d}$, an extended NL prompt $S_{d}$, and an extended tag sequence $T_{d}$—the model requires at least three passes. For each dataset entry, we randomly select one of the seven task types to apply and randomly decide how to split the simple prompt from the detailed prompt ($p_{s}$ from $p_{d}$) and the simple tag set from the detailed tag set ($T_{s}$ from $T_{d}$).
This randomization in task selection and prompt splitting effectively increases the training dataset’s size, allowing the model to encounter a broader range of inputs and outputs. This approach enhances the model’s ability to generalize across the various tasks defined in TIPO.
(a) Direct LLM
(b) Prompt DB
(c) RL (Promptist)
(d) TIPO (Ours)
Figure 4: Scenery test examples from 4 different methods.
6 Evaluation
In this section, we detail the experimental setup used to evaluate the TIPO framework. We outline the formatting conventions for prompts and metadata, define the various tasks employed during training, and describe the models and datasets utilized in our experiments.
6.1 Experimental Setup
To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance): Traditionally, the Frechet Inception Distance (FID) [24] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [44]. Therefore, we measure the Frechet Distance using DinoV2 [35] outputs. In our experiments, we used the pretrained ViT-g [18] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score: We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [17]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score: The AI Corrupt Score is obtained from the AICorruptMetrics [32] in sdeval [33]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
All results presented in this section are tested on the TIPO-200M model.
6.2 Scenery Tag Test
In this test, we randomly select 32,768 entries from Danbooru2023 [50, 34] that include the “scenery” tag. We set up the following inputs:
  • “Scenery” + Meta:
    • Retain all metadata categories as described earlier.
    • Only include “scenery” as the content tag.
  • “Scenery” + Meta + TIPO:
    • Use “scenery” + meta as input.
    • Extend $T_{s}$ (scenery) to $T_{d}$ and generate $p_{d}$ from $T_{d}$.
For the scenery tag test, we generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [37] model on the Danbooru dataset [34].
Figure 4 illustrates the comparative visual result by various methods. The results reveal significant biases in Prompt DB-based [21] and RL-based [22] approaches. The former consistently generates snowy mountain scenes, while the latter produces near-future fantasy scenery, often resulting in corrupted images. The Direct LLM method, utilizing GPT-4o [2], shows improved results but still exhibits notable bias; out of 9 generated prompts, 8 exclusively describe natural scenes only without any human-made structures.
In contrast, TIPO demonstrates superior performance in reproducing the dataset’s distribution. While primarily generating nature scenes, it also incorporates architectural elements and artifacts, more accurately reflecting the diverse content of the original dataset. This balanced output suggests that TIPO effectively mitigates the biases observed in other methods.
Metric Original LLM Prompt DB Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Table 1: Metric results on Scenery Tag Test.
As shown in Table 1, applying TIPO significantly improves FDD performance and AI Corrupt score. The improvement in FDD score suggests that TIPO-generated prompts lead to images that more closely align with the original dataset distribution. Furthermore, the enhanced AI Corrupt score indicates that TIPO aids in generating more coherent images, potentially reducing corruption or distortion in the output.
The results from the direct LLM approach (GPT4o-mini) exemplify a key concern in prompt engineering: while achieving high aesthetic scores, it performs poorly in terms of FDD. This discrepancy underscores the potential pitfall of using LLMs for prompt generation without considering dataset alignment. Such models may produce well-constructed prompts that yield visually appealing images but fail to capture the nuanced distribution of the text-to-image dataset. In contrast, TIPO demonstrates a balance between aesthetic quality and dataset fidelity, offering a more robust solution for prompt optimization in text-to-image generation tasks.
6.3 Short/Truncated Long Test
This test evaluates TIPO’s performance on enhancing natural language prompts which are already contains some useful info for model to generate reasonable images. We randomly selected 10k entries each from [9] and [25]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts for our test. Using these prompts, we generated one image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [37].
Short Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Truncated Long Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Table 2: Metric result in Short/Truncated-long test of prompt generation
Table 2 presents the results of our Short/Truncated-long test for prompt generation. Unlike the scenery tag test (Section 6.2), we observe that the original prompts in this test already produce decent images. Interestingly, all four methods result in worse FDD scores compared to the original prompts, suggesting that the prompt generators introduced new elements not present in the selected 10,000 entries. Given the size of the training set, this outcome is not entirely unexpected.
Examining the Aesthetic and AI Corrupt Scores, we notice that the Direct LLM approach achieves the highest aesthetic score, mirroring the results from the scenery tag test. However, its AI Corrupt Score and FDD value indicate potential issues with OOD inputs, which may occasionally hinder the model’s ability to generate coherent images from its prompts.
TIPO, on the other hand, demonstrates a more balanced performance. It achieves the best AI Corrupt Score, the second-best Aesthetic Score, and a reasonable FDD value. This balanced profile suggests that TIPO is the most suitable choice for real-world applications of modern text-to-image models, where a combination of aesthetic appeal, image coherence, and distribution alignment is crucial.
7 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. By leveraging Large Language Models (LLMs) to extend and refine user-provided prompts, TIPO effectively bridges the gap between simple user inputs and the detailed, content-rich prompts required for high-quality image generation.
Our comprehensive evaluation demonstrates TIPO’s effectiveness across multiple dimensions of image quality and fidelity. The framework consistently produced prompts that led to images closely aligned with original dataset distributions, while simultaneously improving aesthetic appeal and reducing the likelihood of corrupted or nonsensical outputs. This balanced performance underscores TIPO’s robustness and suitability for real-world applications of T2I models, where a combination of distribution alignment, visual appeal, and image coherence is crucial.
Notably, our experiments highlight the critical role of prompt engineering in advancing T2I technology. The improvements achieved through TIPO’s prompt modifications often rivaled or surpassed performance differences between various T2I model architectures. This finding emphasizes that effective prompt engineering can significantly enhance model performance without necessitating architectural changes, potentially offering a more efficient path to improving T2I systems.
The success of TIPO opens several promising avenues for future research:
  1. 1.
    Broader Application: Exploring TIPO’s applicability to other generative tasks, such as text generation or audio synthesis, could further demonstrate the versatility and impact of automatic prompt engineering.
  2. 2.
    Interactive Systems: Incorporating TIPO into interactive applications where users can iteratively refine prompts in real-time may enhance user experience and enable more precise control over generated content.
  3. 3.
    Advanced Alignment Techniques: Investigating more sophisticated methods for aligning user inputs with dataset distributions could further enhance the performance and reliability of generative models.
  4. 4.
    Cross-Modal Applications: Adapting TIPO’s principles to facilitate cross-modal generation tasks, such as text-to-video or image-to-text, could expand its utility in multimodal AI systems.
In conclusion, TIPO represents a significant advancement in the field of text-to-image generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts, TIPO not only improves the quality and specificity of generated images but also highlights the importance of aligning user inputs with underlying data distributions. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in AI-generated content.
Appendix
Appendix A Dataset/Resource
A.1 Danbooru2023
The Danbooru2023 dataset [50, 49, 34] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with highly specific and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing a wide variety of visual content with granular annotations. These annotations allow for the creation of nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
Key Characteristics:
  • Rich Annotations: Danbooru2023’s detailed tags enable the differentiation of subtle variations in image content, such as “long hair” versus “very long hair,” which is crucial for the specificity required in high-quality image generation.
  • Large Volume: The dataset’s extensive size provides a robust foundation for training models, ensuring that they can learn from a diverse array of images and annotations.
  • Tag-Based Prompting: Utilizing the detailed tags from Danbooru2023, the LLM can generate refined prompts that lead to more accurate and high-quality image generation by the T2I model.
A.2 GBC10M
The GBC10M dataset [25] is a large-scale collection of 10 million images sourced from CC12M [10], annotated using the Graph-Based Captioning (GBC) approach. Each image in GBC10M is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define the hierarchical relationships among them. Annotations are generated automatically through a pipeline that leverages pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing both detailed descriptions and structural information, improving model performance in downstream tasks. All data is provided in JSON lines format, containing image URLs, bounding boxes, and caption annotations.
In TIPO, we only utilize the root node from GBC10M, since it provides both detail and short captions.
A.3 Coyo HD 11M
The Coyo HD 11M dataset [9] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The ”HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [30] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
Appendix B Implementation Detail
In this appendix, we provide all the missing details about our dataset construction process that are not mentioned in Section 4.4 and 5.2.
B.1 Training Settings
TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
Architecture LLaMA
Type Pretrain Finetune Pretrain
Vocab Size 32013 - 32013
Hidden Dim 768 - 1280
Attention Heads 12 - 20
MLP Dim 2304 - 3840
Hidden Layers 20 - 20
Model Parameters 203M - 508M
Max Learning Rate 2e-4 5e-5 2e-4
Optimizer AdamW AdamW AdamW
betas 0.9, 0.98 0.9, 0.98 0.9, 0.98
weight decay 0.01 0.01 0.01
Dataset GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
max context length 512 1024 1024
global batch size 2048 2048 3584
Token Seen 22.625B 18.339B 31.274B
Hardware 4 x RTX3090 4 x RTX3090 8 x H100
Training Time (wall) 150 hour 270 hour 100 hour
Table 3: Training settings and details for TIPO-200M and TIPO-500M.
B.2 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts. Figure 5 illustrates this comprehensive workflow.
Figure 5: The inference workflow in TIPO, demonstrating how different task combinations are used to refine or expand tag-based and natural language prompts.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
Appendix C Evaluation Statistics
In this appendix, we provide more statistics for the result obtained in Section 6.
C.1 Scenery Tag Test
(a) The box plot for the Aesthetic Score result of scenery tag test.
(b) The box plot for the AI Corrupt Score result of scenery tag test.
Figure 6: The box plot for scenery tag test.
The box plot shown in Figure 6 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 6.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
C.2 Short/Truncated Long Test
Figures 7 and 8 display the box plots of aesthetic scores and AI corruption scores obtained from the short/truncated long test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
(a) The box plot for the Aesthetic Score result of short caption input.
(b) The box plot for the AI Corrupt score result of short caption input.
Figure 7: The box plot for short caption input.
(a) The box plot for the Aesthetic Score result of truncated long caption input.
(b) The box plot for the AI Corrupt score result of truncated long caption input.
Figure 8: The box plot for truncated long caption input.
Appendix D TIPO example
In this appendix, we provide some text example of TIPO’s input and output.
TIPO Format template User Input:
1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo
Formatted TIPO Output after Expand Tags and Expand Natural Prompt:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting.
Formatted Output For Text-to-Image
1girl, ciloranko, maccha \(mochancc\), ningen mame, ask \(askzy\), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe
Refer to caption User InputRefer to caption TIPO Output
Figure 9: An example formatted content we used for training and inference in TIPO.
TIPO Format template User Input:
scenery, no humans, masterpiece, absurdres, newest, safe
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> tag: scenery, no humans
Formatted TIPO Output after Expand Tags and tag_to_long task:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> <|tag_to_long|> tag: scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar (cafe), lamp, road long: A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. on top of the traffic light pole, there seems to be an air conditioning unit. the street is lined with trees and bushes, and there is graffiti on the ground.
Formatted Output For Text-to-Image
scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar \(cafe\), lamp, road, A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. masterpiece, newest, absurdres, safe
Refer to caption User InputRefer to caption TIPO Output
Figure 10: An example formatted content we used for training and inference in TIPO.
Appendix E Image Examples
In this section, we present sample images from the experiments described in Section 6 to visually demonstrate the improvements achieved by TIPO.
E.1 Scenery Tag Test
(a) Scenery Tag Only
(b) TIPO + Scenery Tag
Figure 11: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
Figure 11 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [50]. The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
E.2 Short/Truncated Long Test
(a) Short Caption
(b) TIPO-Generated Caption
(c) Truncated Long Caption
(d) TIPO-Extended Caption
Figure 12: Comparison of generated images using original input (left) vs. TIPO-enhanced input (right)
Figure 12 illustrates the differences between short captions, truncated long captions, TIPO-generated captions, and TIPO-extended captions. The “short prompt” and “truncated long prompt” used in this experiment typically consist of 1-2 sentences, resulting in reasonably good quality outputs. However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
Appendix F Topic Distribution Visualization
To analyze the meaningfulness and diversity of topics within text inference results, we implemented a multi-stage topic modeling and clustering methodology. Initially, we employed Latent Dirichlet Allocation (LDA) to extract varying numbers of topics—specifically 20, 30, 50, and 100—from the corpus [8]. This approach aimed to identify representative words that are significant within the dataset while filtering out common stop words, ensuring that nonsensical or irrelevant terms are excluded from topic classification.
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics (Stevens et al., 2012).
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [23], shown in Table 4, 5, and 6. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Table 4: Inertia for COYO-Dataset inference, higher is better
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 170.78 137.94 1078.38 204.71 125.90 144.59 1037.44 271.77
30 461.79 417.74 1758.40 736.74 327.48 195.13 1323.88 512.07
50 829.68 730.73 861.15 1036.33 400.90 373.74 823.51 959.18
100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
Table 5: Inertia for GBC-Dataset inference, higher is better
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 184.53 352.16 244.79 246.15 125.47 198.94 278.56 211.98
30 452.01 566.74 505.56 441.34 204.28 328.70 372.07 471.28
50 571.77 895.47 1227.30 990.17 438.89 313.48 737.65 788.41
100 1291.60 1742.36 1675.41 1550.32 631.61 628.78 1573.47 1855.90
Table 6: Inertia for Scenery extend inference, higher is better
Size PromptDB GPT4o-mini Promptist TIPO
20 60.82 734.52 210.60 139.29
30 275.76 1141.77 415.95 355.20
50 630.50 826.29 722.75 1002.36
100 2026.39 1879.08 802.93 1883.70
We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
Figure 13: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.