### v2

# TIPO: Text to Image with Text Presampling for Prompt Optimization

Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation. Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [47, 7, 21, 55, 51, 52, 57, 53, 48, 56, 15, 14, 34, 22, 8]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering.
Previous works have attempted to extend prompts with optimization through various methods. Some approaches use LLMs to expand prompts [41]; however, due to the misalignment between the LLM’s text-rich training corpus and the T2I model’s caption-focused corpus, only a limited set of manually-engineered prompts yield effective results. Other methods rely on curated prompt sets [5, 17], which, while potentially incorporating user preferences, fail to fully leverage the T2I model’s trained prompt distribution. Additionally, reinforcement learning-based techniques [26] require multi-turn inference, incurring high computational costs, and often struggle to generate diverse outputs due to the model’s inherent corpus limitations.
To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
Our work makes the following key contributions:
  1. 1.
    We introduce an universal prompt optimization framework that is designed to work with any text-to-image model, providing a versatile solution for improving image generation across different architectures.
  2. 2.
    We provide comprehensive evaluation methodology and results including using both standard and non-standard metrics. Our experiments demonstrate that TIPO’s prompt engineering method can improve image quality in terms of mathematical measures and user preferences, validating its effectiveness from multiple perspectives.
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
This paper presents TIPO’s theoretical foundations (Section 3), implementation details (Section 4), and experimental results (Section 5) demonstrating its effectiveness in improving aesthetic scores [19], reducing image corruption [43, 44], and better aligning generated images with dataset distributions [28, 46]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
2 Related Works
2.1 Direct Prompt Refinement Through LLMs
With the advent of advanced LLMs, several projects have leveraged their few-shot in-context learning capabilities to optimize prompts for Text-to-Image (T2I) models. For instance, CogView3 [67] and the work by Lee et al. [32] utilize GPT-J and Text Style Transfer (TST) techniques for prompt refinement. However, this approach relies heavily on the LLM’s inherent knowledge of visual content descriptions, which can be unpredictable and may not always align with specific T2I model requirements.
2.2 Extending Prompts with Manually Collected Data
In the open-source T2I community, a common approach involves using curated datasets of “good prompts” to fine-tune or train LLMs for prompt extension [5, 60, 17]. These models are typically trained on image-text pairs collected from users of proprietary T2I services like Midjourney [^1] https://www.midjourney.com or Dall-E 3 [7]. While this method is straightforward and somewhat effective, it often lacks the versatility to adapt to diverse T2I models beyond the one used in the original dataset collection.
2.3 Image Output-Based Tuning
Recent works have explored more sophisticated prompt optimization techniques based on image outputs.
Reinforcement Learning Approaches: Promptist [26] introduced a two-stage framework combining supervised fine-tuning with reinforcement learning (RL). The RL stage incorporates aesthetic and relevance scores as environmental feedback, enabling the model to learn optimal prompt refinement strategies.
Iterative Optimization: OPT2I [41] proposed a backpropagation-free prompt optimization framework. This method uses scorers and LLMs to iteratively update input prompts, aiming to improve aesthetic quality and coherence.
These RL-based or RL-like approaches allow for prompt refinement tailored to specific model properties or user preferences. However, they often require extensive computational resources and may not generalize well across different T2I models or datasets.
While these methods have shown promise, they often face challenges such as difficulty in maintaining alignment with original user intent, challenges in collecting manually engineered prompts, high computational costs for RL-based methods, and limited generalization across diverse T2I models. By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
[…]
3 Prompt Optimization
3.1 Motivation and Challenges
Nowadays text-conditioned models [8, 48, 21] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the familiar vocabulary and stylistic tendencies of the model [25], while others rely on direct feedback from the model to iteratively improve prompt alignment [26]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text input. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
3.2 Formulation
To formalize prompt extension, we define $\mathcal{P}$ as the set of all possible prompts, $\mathcal{N}$ as the Gaussian noise vector space, and $\mathcal{I}$ as the set of all possible images. A T2I model then acts as a mapping function:
$f(p):\mathcal{N}\rightarrow\mathcal{I}_{p}$
where, for any prompt $p\in\mathcal{P}$, the model maps noise vectors from $\mathcal{N}$ to a subset of images $\mathcal{I}_{p}\subseteq\mathcal{I}$ associated with the prompt $p$.
Now we define a pair of prompts: simple prompt $p_{s}$ and detailed prompt $p_{d}$. Formally, for a simple prompt $p_{s}$ and a detailed prompt $p_{d}$ that extends $p_{s}$, we have the following relationship:
$f(p_{d})(\mathcal{N})\subseteq f(p_{s})(\mathcal{N})$
This formulation implies that the image subset generated by the detailed prompt $p_{d}$ is contained within the subset generated by the simple prompt $p_{s}$, thereby preserving the relevance of the initial prompt while adding specificity.
To further encapsulate the prompt generation process, we define the prompt extension function, $\mathcal{E}$. This function takes a simple prompt $p_{s}$ and additional metadata $\mathcal{M}$ (such as tags) as inputs, and produces a detailed prompt $p_{d}$ as output:
$\mathcal{E}:\mathcal{P}_{s}\times\mathcal{M}\rightarrow\mathcal{P}_{d}$
where $\mathcal{P}_{s}$ is the space of simple prompts and $\mathcal{P}_{d}$ is the space of detailed prompts. Here, $\mathcal{M}$ provides auxiliary information that guides the extension of $p_{s}$ into a more refined prompt $p_{d}$.
For each instance of prompt extension, we apply the function as follows:
$p_{d}=\mathcal{E}(p_{s},M)$
where $M$ could represent a specific metadata component, such as $T_{s}$, or other forms of contextual information depending on the input type.
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images or the model to generate not user intent images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory, as shown in Figure 2 (c). For example, the generated images may deviate from the user’s intent by not depicting an astronaut (left top), producing a horse-shaped tank (right top) instead of a horse, omitting Mars (left bottom), or even including a random snake (right bottom), which introduces nonsensical elements. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\supseteq\mathcal{I}\supseteq\mathcal{I}_{p_{s}}$
Which means:
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
This scenario is illustrated in Figure 2 (c), where “Simple Prompt + Random Words” shows a pointy distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework and Model
4.1 Distribution-Aligned Prompt Extension
Central to TIPO’s method is an extension function $\mathcal{E}$, which iteratively refines prompts to enhance both the diversity and quality of generated images while remaining aligned with the intent of the original user-provided prompt and the training data distribution. By applying $\mathcal{E}$ to a simple prompt $p_{s}$ to generate a detailed prompt $p_{d}$, TIPO aims to approximate the target distribution:
$\bigcup_{i=0}^{n}f(p_{t_{i}})(\mathcal{N})\approx\mathcal{I}_{p_{s}}$
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 (d) “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
4.2 Prompt Structure
The construction of simple prompts, $p_{s}$, and their corresponding detailed prompts, $p_{d}$, varies depending on the dataset and input modality. We consider two primary scenarios: tag-based captions and natural language (NL) captions.
Tags are used to describe the content of images and are commonly formatted as a comma-separated string when input into T2I models. For a given image-tag pair, let $T_{n}=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ represent the complete set of tags for an image. We derive the simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ where $m<n$, and define the detailed tag set as $T_{d}=T_{n}$. The simple and detailed tag prompts are then constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{d}=\text{concat}(T_{d})$
NL captions are typically generated by LMs with vision modality [36, 1, 2, 33, 6, 16, 63, 18] and may contain multiple sentences. Each image is associated with either two types of NL captions—a short NL caption $S_{s}$ providing a brief, general description, and a long NL caption $S_{d}$ offering an extended, detailed description—or a single long caption that can be segmented. For images with a single long caption, let
$S_{d}=\{\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}\}$
represent the complete set of sentences in $p_{d}$. We derive the simple sentence set
$S_{s}=\{\text{sentence}_{1},\text{sentence}_{2},\ldots,\text{sentence}_{m}\}$
where $m<n$, and define the detailed sentence set as $S_{d}=S_{n}$. The simple and detailed NL prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{d}=\text{concat}(S_{d})$
These ensure that $p_{d}$ is a superset of $p_{s}$, aligning with the concept that detailed prompts provide a more specific description.
To generate $p_{d}$ from $p_{s}$, TIPO employs two distinct formats based on the relationship between $p_{s}$ and $p_{d}$:
  1. 1.
    When $p_{s}$ is a substring of $p_{d}$: The format <meta> <p_d> is used. As TIPO model is an autoregressive LM, $p_{d}$ is directly used after the <meta> tokens, allowing the model to generate subsequent tokens based on any substring of $p_{d}$.
  2. 2.
    When $p_{s}$ is not a substring of $p_{d}$: The format <meta> <p_s> <p_d> is employed. Since $p_{s}$ is not contained within $p_{d}$, both are included separately after the <meta> tokens to guide the generation process effectively.
4.3 Prompt Formatting
In TIPO, we define a simple and consistent format for the content of the <meta> token or the simple prompt $p_{s}$. The format is as follows: <Category>: <Content>. Metadata categories might include artist, copyright, aspect ratio, quality, and year. For instance,
quality: masterpiece, year: newest
Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
4.4 Task Definitions
Figure 3: TIPO’s task flow within a single generation cycle. Starting from any node, each arrow represents a sequential extension step, with other prompt used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
TIPO defines three core tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. Each task is represented by an arrow in Figure 3, illustrating how TIPO can use tag and NL prompts interchangeably as input sources to generate detailed outputs within a single generation cycle. When an arrow cycle is established, TIPO can complete multiple extension within a single generation pass.
For example, TIPO can start from a short NL prompt to produce a detailed tag sequence, or it can combine tasks like short_to_tag and tag_to_long to create a composite task, short_to_tag_to_long, which generates an extended NL prompt from an initial short NL input via tag generation. Further details on specific task sequences are provided in Appendix.
4.5 An Example Scenario and Workflow
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence. During each generation cycle, one input type is treated as the primary prompt which is target prompt for extending or refining, while the other types are used as <Content> of metadata to enrich the context. Figure 4 provides an example of this setup, where:
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\end{array}$
The generation proceeds sequentially as follows:
  1. 1.
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$
  2. 2.
    tag_to_long: $T_{d}$ is incorporated into the metadata, and TIPO produces a refined short NL prompt $S_{s}$ based on $T_{d}$.
  3. 3.
    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
  4. 4.
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context rich prompt, $p_{d}$.
This iterative process allows TIPO to build prompts that are both detailed and contextually aligned with user inputs, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
4.6 Model Architecture
TIPO utilizes a causal autoregressive language model framework as its prompt extension function, $\mathcal{E}$, to iteratively extend and refine user-provided prompts into more detailed, content-rich prompts $p_{d}$ based on simpler inputs $p_{s}$. Our implementation leverages the LLaMA architecture [61, 62, 3], employing models with 200M and 500M parameters.
  • TIPO-200M: Pretrained on the Danbooru2023 [45, 65] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
Throughout training, non-padding tokens were used for token count calculations, with the effective count reflecting the relatively short and variable length of data entries.
4.7 Training Process
In each training pass, TIPO applies dataset augmentation by selecting tasks from all possible task combinations such as short_to_tag_to_long, including extending tag sequences, NL prompt extension, and refined NL prompt generation. This random selection across all task combinations significantly broadens the dataset’s effective scope, enhancing model generalization without additional manual engineering or runtime tuning.
(a) Direct LLM
(b) Prompt DB
(c) RL (Promptist)
(d) TIPO (Ours)
Figure 5: In-domain test regard to scenery test examples from 4 different methods.
(a) Scenery Tag Only
(b) Scenery Tag + TIPO
(c) Truncated Long Prompt
(d) Truncated Prompt + TIPO
Figure 6: Comparison of generated images: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated long prompt, and (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images even from minimal prompts.
5 Evaluation
5.1 Experimental Setup
In this section, we detail the experimental setup used to evaluate prompt optimization framework. To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance).
Traditionally, the Frechet Inception Distance (FID) [28] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [58]. Therefore, we measure the Frechet Distance using DinoV2 [46] outputs. In our experiments, we used the pretrained ViT-g [20] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score.
We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [19]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score.
The AI Corrupt Score is obtained from the AICorruptMetrics [43] in sdeval [44]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
Vendi Score.
To quantify the diversity of generated images, we employ the Vendi Score metric based on DinoV2 [46] embeddings. The computation involves calculating a similarity matrix $K$ using cosine similarity between image embeddings, followed by computing the von Neumann entropy of the normalized matrix $K/n$, adhering to the standard Vendi Score methodology. This metric provides a quantitative measure of the distributional diversity within a set of generated images. Higher Vendi Scores indicate greater inter-image diversity, reflecting the model’s capability to generate varied and distinct outputs.
Extension Functions.
We used both GPT-4o-mini [2] and LLaMA-3.2-1B [3] as Direct LLM models for prompt generation, balancing performance and efficiency. GPT-4o-mini, as a lightweight closed-source model, offers speed and affordability, while LLaMA-3.2-1B represents a similar scale to TIPO in open-source models. Also, Prompt DB [17] and Promptist [26] serve as baseline methods for prompt extension.
5.2 In-domain test regard to scenery tag
To evaluate the model’s capability to generate diverse and accurate outputs, we chose the scenery tag, which includes both natural landscapes and human-made structures. In this test, we randomly select 32,768 entries from Danbooru2023 [65, 45] that include the “scenery” tag. We set up the following prompt:
<meta> + scenery + $\mathcal{E}$>
Then, we apply extension functions to extend only the scenery, $T_{s}$, as the <Content> tag. We generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [48] model on the Danbooru dataset [45].
Metric Original LLM Prompt DB Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Table 1: Metric results on in-domain test. Bold values indicate the best performance, and underlined values indicate the second-best. Original refers to images generated using the original text captions; LLM uses GPT-4o-mini [2]; the remaining columns represent results from Prompt DB [25], Promptist [26], and TIPO, respectively.
As shown in Table 1, TIPO-generated prompts improve FDD scores by aligning images more closely with the original dataset distribution. Enhanced AI Corrupt scores show that TIPO reduces image corruption, creating more coherent outputs by replacing vague descriptions (e.g., detailed background) with specific elements (e.g., garden filled with forget-me-nots and blue butterflies). The direct LLM approach highlights a key issue in prompt engineering: it achieves high aesthetic scores but compromises FDD, reducing the diversity of images.
We provide more comparative visual result in Figure 5. The results reveal significant biases in Prompt DB  [25] and RL-based [26] approaches. PromptDB and Direct-LLM method exhibits notable bias, by generating mountain/lake scene majorly. Meanwhile, Promptist produces diverse near-future fantasy sceneries, however often produces corrupted images due to unreasonable vocabulary usage (4K, masterpiece, instagram), supposedly introduced by user prompts.
By contrast, TIPO consistently reproduces a balanced output distribution, incorporating both natural and architectural elements, effectively mitigating the biases observed in other methods.
(a) Original prompt best aesthetic
(b) GPT-4o-mini best aesthetic
(c) TIPO best aesthetic
Figure 7: Top-10% aesthetic score images generated by the SD-3.5-Large model [19]. Rows show samples from Original, Direct LLM-Modified prompts, and TIPO-extended prompts. Direct LLM images focus primarily on natural scenery, while generated image from TIPO-extended prompt display a wider variety, including urban scenes and human figures.
5.3 In-domain prompt generation test
We aim to test each models’ performance on regard to reconstructing / upsampling the given prompt, in given dataset distribution. By comparing original ground truth and generated outputs, we can directly compare FDD, aesthetic score, AI corrupt scores.
We randomly selected 10k entries each from [12] and [29]. For each entry, we extracted the short caption and the truncated long caption (limited to a maximum of two sentences). These captions serve as user input prompts $p_{s}$ for in-domain prompt generation test. Using these prompts, we generated single image per prompt with the SDXL-1.0-base [^3] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0 model [48].
Short Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Truncated Long Original LLM Prompt DB Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Table 2: Performance comparison in the in-domain prompt generation test. Bold values indicate the best performance, and underlined values indicate the second-best. Direct LLM achieves the highest Aesthetic Score but struggles with OOD issues. TIPO provides a balanced outcome, excelling in AI Corrupt Score and maintaining strong alignment, making it ideal for practical applications.
Table 2 presents the results of our Short/Truncated caption test for prompt generation. Compared to in-domain test, Section 5.2, which might be considered slightly out of domain by SDXL model, we observed that prompts covered by this test produces meaningful, coherent images with sufficient detail level. FDD scores were degraded compared to original-ground truth prompts, not truncated long caption-which implies prompt extension functions resulted lack of diversity, or unmatched components in generation.
We observe that Direct LLM approach achieves the highest aesthetic scores by wide vocabulary with user preference matching, however results low FDD, AI corruption score, which implies unmatched vocabularies are being used in-prompt.
Original LLM Prompt DB Promptist TIPO(Ours)
Top 10% $\uparrow$ 10.98 7.84 12.17 10.12 17.33
Worst 10% $\uparrow$ 12.10 9.76 15.23 14.14 13.92
Table 3: Vendi Scores [23] for the top and worst 10% aesthetic images across prompt generation methods. Higher value indicate higher diversity of samples. TIPO achieves high score in the top 10% (17.33), showing greater visual diversity as well as worst 10%. Result suggests TIPO’s balanced and robust variety in generated images. Meanwhile, LLM has the lowest scores in both categories, indicating limited diversity and repetition.
5.4 Out of focus evaluation.
In this section, we analyze TIPO’s behavior in out-of-focus model, SD-3.5-Large [^4] https://huggingface.co/stabilityai/stable-diffusion-3.5-large built upon MMDiT architecture [21]. We compose original tag, natural language prompt created by GPT-4o-mini, tag+natural language extended prompt by TIPO (ours), then feed to SD-3.5-Large inference pipeline.
Analysis of Vendi Scores [23] (Table 3) reveals contrasting patterns: Direct LLM-generated images show limited diversity despite high aesthetic scores, while TIPO maintains conceptual diversity across all quality levels. However, TIPO’s training on diverse datasets [45, 65, 12, 29] introduces domain-specific elements that may not align with SD3.5L’s training distribution. This mismatch demonstrates our core finding: prompt generation models perform optimally when trained on domain-matched prompt corpora.
The balanced profile of TIPO suggests that TIPO can work as suitable solution for prompt upsampling task, which can accurately project trained prompt with few initial word input.
6 Conclusion
In this paper, we introduced TIPO (Text to Image with text Presampling for Prompt Optimization), a novel framework designed to enhance the quality of Text-to-Image (T2I) models through automatic prompt engineering. TIPO represents a significant advancement in the field of T2I generation, offering a versatile and efficient solution for prompt optimization. By automating the refinement and extension of user prompts. We also highlights the importance of aligning user inputs with underlying data distributions through the comparisons between different methods for prompt engineering. As generative models continue to evolve, frameworks like TIPO will be instrumental in unlocking their full potential, ensuring they meet diverse user needs with precision and creativity while pushing the boundaries of what’s possible in generative models.
Appendix
[…]
Appendix A Dataset/Resource
A.1 Danbooru2023
The Danbooru2023 dataset [65, 64, 45] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with highly specific and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing a wide variety of visual content with granular annotations. These annotations allow for the creation of nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
Key Characteristics:
  • Rich Annotations: Danbooru2023’s detailed tags enable the differentiation of subtle variations in image content, such as “long hair” versus “very long hair,” which is crucial for the specificity required in high-quality image generation.
  • Large Volume: The dataset’s extensive size provides a robust foundation for training models, ensuring that they can learn from a diverse array of images and annotations.
  • Tag-Based Prompting: Utilizing the detailed tags from Danbooru2023, the LLM can generate refined prompts that lead to more accurate and high-quality image generation by the T2I model.
A.2 GBC10M
The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach. Each image in GBC10M is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define the hierarchical relationships among them. Annotations are generated automatically through a pipeline that leverages pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing both detailed descriptions and structural information, improving model performance in downstream tasks. All data is provided in JSON lines format, containing image URLs, bounding boxes, and caption annotations.
In TIPO, we only utilize the root node from GBC10M, since it provides both detail and short captions.
A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-density, high-definition images, with 22.8 million synthetic captions generated from the Coyo-700M dataset. The “HD” refers to both the high resolution of images (filtered to have at least 512 pixels on the shortest edge) and the high concept density, addressing the issues found in previous alt-text image pair datasets, which often included low-quality or low-concept images. Synthetic captions for this dataset were produced using the LLaVA-Next-8B model [37] built upon LLaMA 3 [3] with post-processing for conciseness and clarity. This dataset provides high-quality visual and textual pairs, ideal for training advanced vision-language models.
In TIPO, we utilize the short/long captions, booru tags, and open images tags from Coyo HD 11M.
A.4 Stable Diffusion XL
The Stable Diffusion XL [48] improves upon earlier versions [54] with a significantly larger UNet backbone and dual text encoders, CLIP ViT-L [49] and OpenCLIP ViT-bigG [30], to enhance text conditioning. It supports resolutions up to 1024×1024 and is designed for diverse image generation. Although explicit dataset details are not provided, figures in the paper confirm that SDXL accepts both natural language (NL) prompts and tags as input.
A.5 Stable Diffusion 3.5
The Stable Diffusion 3.5 (SD-3.5) incorporates a cutting-edge architecture, MMDiT [21], along with the Rectified Flow [38, 4, 35] formulation for improved text-to-image generation. The model utilizes triple text encoders—CLIP/ViT-L, OpenCLIP/ViT-G, and T5-XXL [50]—for robust text conditioning. SD-3.5 supports resolutions up to 1024×1024 and employs a 50/50 mix of original and CogVLM-generated captions for training. While explicit dataset details are not provided, figures in the paper confirm its capability to process both natural language prompts and tags.
In this study, we use the SD-3.5-Large [^8] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s property which are not mentioned in Section 4.3 and 4.4.
B.1 Training Data Construction
This section details our methodology for constructing and preprocessing training data to ensure robust model performance across various input scenarios.
Length Control
To systematically control output prompt length, we implement a structured length categorization system using special length tags. These tags enforce specific constraints on both tag counts and natural language sentence lengths. For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description. We define four distinct length categories, each with strict bounds for both tag count and sentence length.
Type Very Short Short Long Very Long
Tags (count) 18 36 48 72
NL (sentences) 2 4 8 18
Table 4: Maximum length specifications for each category and caption type. For each category, the actual count/length must not exceed these values.
Random Augmentation
To enhance input diversity and better simulate real-world usage patterns, we implement several data augmentation strategies:
  • Metadata Tags: For tags representing image metadata (e.g., artist, character, aspect ratio), we employ two randomization techniques:
    • Random removal of metadata tags
    • Random repositioning of metadata tags to the end of the prompt, after all content-related descriptions
    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from the content descriptions.
  • Content Tags: For tags describing image content (e.g., objects, actions, attributes), we implement:
    • Random shuffling of tag order within the content section
    • Length-based truncation to meet target length constraints while preserving key content information
  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve both context-setting opening sentences and concluding details. This maintains coherent narrative flow while meeting target length requirements.
These augmentation strategies create a more diverse training dataset that better reflects real-world prompt variations, improving the model’s robustness and adaptability to different input styles and formats.
B.2 Training Settings and Model Configurations
TIPO-100M TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
Architecture LLaMA
Type Pretrain Pretrain Finetune Pretrain
Vocab Size 32013
Hidden Dim 640 768 - 1280
Attention Heads 10 12 - 20
MLP Dim 2240 2304 - 3840
Hidden Layers 10 20 - 20
Model Parameters 100M 203M - 508M
Max Learning Rate 5e-4 2e-4 5e-5 2e-4
Optimizer AdamW
LR scheduler Cosine Annealing LR
betas 0.9, 0.98
weight decay 0.01
Dataset Coyo, GBC, Dan GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
Epoch 1 5 3 5
max context length 512 512 1024 1024
global batch size 1024 2048 2048 3584
Token Seen 6.0240B 22.625B 18.339B 31.274B
Hardware 4 x RTX3090 4 x RTX3090 4 x RTX3090 8 x H100
Training Time (wall) 22.5 hour 150 hour 270 hour 100 hour
Table 5: Training settings and details for TIPO-100M, 200M, and 500M. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
Tokenizer and Task Tokens
TIPO employs a vocabulary derived from LLaMA2 [62] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders. This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
  • Placeholder Token (1 token):
        <|empty|>
  • Task Tokens (8 tokens):
        <|gen_meta|>, <|tag_to_long|>, <|short_to_tag|>,
        <|long_to_tag|>, <|short_to_long|>, <|short_to_tag_to_long|>,
        <|short_to_long_to_tag|>, <|tag_to_short_to_long|>
  • Length Tokens (4 tokens):
        <|very_short|>, <|short|>, <|long|>, <|very_long|>
Optimizer and Learning Schedule
Training is performed using the AdamW optimizer [39], with a cosine annealing learning rate scheduler [40]. The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. Maximum learning rates are adjusted per model size, as outlined in Table 5.
Training Configurations
TIPO models are trained in multiple stages. Table 5 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and finetuning was conducted on datasets like Danbooru2023 [45], GBC10M [29], and CoyoHD11M [12].
Augmented Task Representation
Each dataset entry undergoes random task assignment and splitting to simulate a wide range of input-output mappings, effectively increasing the dataset size. For example, a single entry may contribute to tasks like short_to_tag or tag_to_long, with length modifiers dynamically controlling the output verbosity. This approach ensures the model can handle diverse tasks while maintaining robust generalization.
Hardware and Time Requirements
Training was conducted on NVIDIA RTX3090 GPUs for smaller models and H100 GPUs for TIPO-500M. Total wall-clock training times ranged from 22.5 hours for TIPO-100M to 270 hours for fine-tuning TIPO-200M.
Token Seen and Effective Training
Non-padding tokens are used to measure the effective token count during training, ensuring efficiency given the short and variable data lengths. Table 5 details the total tokens seen per model and training stage, illustrating the comprehensive exposure to diverse data entries.
[…]
Figure 8: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle. ¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
B.3 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to achieve refinement or expansion of both tag-based prompts and natural language prompts. Figure 8 illustrates this comprehensive workflow.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that both tag-based and natural language prompts are optimized, resulting in more detailed and effective inputs for text-to-image models.
B.4 Model Speed Comparison
We conducted comprehensive speed tests of our 100M, 200M and 500M parameter models using the inference pipeline described in Section 5.2. Each prompt requires two sequential generation steps. Our primary metric is average tokens generated per second, which reflects real-world task performance rather than theoretical maximum throughput.
The evaluation was performed using llama.cpp [24], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
TIPO-100M TIPO-200M TIPO-500M
Hardware Platform tok/sec gen time tok/sec gen time tok/sec gen time
M1 Max (32 GPU cores) 339.4 0.66 190.0 1.23 119.4 2.02
RTX 3090 558.5 0.42 341.4 0.69 289.8 0.81
RTX 4090 742.9 0.29 454.5 0.51 359.7 0.63
Table 6: Model performance comparison across different hardware platforms. Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time required for a complete two-step prompt optimization process.
Appendix C Evaluation Statistics
In this appendix, we provide more statistics for the result obtained in Section 5.
C.1 In-domain test regard to scenery tag
(a) The box plot for the Aesthetic Score result of scenery tag test.
(b) The box plot for the AI Corrupt Score result of scenery tag test.
(c) The KDE plot for the Aesthetic Score result of scenery tag test.
(d) The KDE plot for the AI Corrupt Score result of scenery tag test.
Figure 9: The distribution of aesthetic and ai corrupt score for scenery tag test.
The box plot and Kernel Density Estimation (KDE) plot shown in Figure 9 presents the aesthetic scores and AI corruption scores for results from the scenery tag test described in Section 5.2. The analysis clearly demonstrates that TIPO significantly outperforms all other methods, with a substantial margin of improvement.
C.2 In-domain prompt generation test
(a) The box plot for the Aesthetic Score result of short prompt input.
(b) The box plot for the AI Corrupt score result of short prompt input.
(c) The KDE plot for the Aesthetic Score result of short prompt input.
(d) The KDE plot for the AI Corrupt Score result of short prompt input.
Figure 10: The distribution of aesthetic and ai corrupt score for short prompt input.
(a) The box plot for the Aesthetic Score result of truncated long prompt input.
(b) The box plot for the AI Corrupt score result of truncated long prompt input.
(c) The KDE plot for the Aesthetic Score result of truncated long prompt input.
(d) The KDE plot for the AI Corrupt Score result of truncated long prompt input.
Figure 11: The distribution of aesthetic and ai corrupt score for truncated long prompt input.
Figures 10 and 11 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal that the differences in performance between various methods are subtle, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images compared to other methods.
C.3 Out of focus evaluation
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by Prompt-DB
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 12: The similarity matrix for the 100 best aesthetic results generated in the SD3.5-Large experiments. Off-diagonal elements of the matrix indicate the similarity between different images. A lower value for an off-diagonal element indicates greater diversity among the generated images.
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by Prompt-DB
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 13: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
Figures 12 and 13 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [22]. A matrix with predominantly lower similarity values (overall brighter appearance) indicates high diversity across all generated images, while higher similarity values (overall darker appearance) suggest consistent but less diverse outputs.
For quantitative diversity metrics, please refer to Table 3 in Section 5.
Appendix D TIPO example
In this appendix, we provide some text example of TIPO’s input and output.
TIPO Format template User Input:
1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo
Formatted TIPO Output after Expand Tags and Expand Natural Prompt:
meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting.
Formatted Output For Text-to-Image
1girl, ciloranko, maccha \(mochancc\), ningen mame, ask \(askzy\), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe
Refer to caption User Input Refer to caption TIPO Output
Figure 14: An example formatted content we used for training and inference in TIPO.
TIPO Format template User Input:
scenery, no humans, masterpiece, absurdres, newest, safe
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> tag: scenery, no humans
Formatted TIPO Output after Expand Tags and tag_to_long task:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> <|tag_to_long|> tag: scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar (cafe), lamp, road long: A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. on top of the traffic light pole, there seems to be an air conditioning unit. the street is lined with trees and bushes, and there is graffiti on the ground.
Formatted Output For Text-to-Image
scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar \(cafe\), lamp, road, A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. masterpiece, newest, absurdres, safe
Refer to caption User Input Refer to caption TIPO Output
Figure 15: An example formatted content we used for training and inference in TIPO.
Appendix E Image Examples
In this section, we present sample images from the experiments described in Section 5 to visually demonstrate the improvements achieved by TIPO.
E.1 In-domain test regard to scenery tag
Figure 16: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
Figure 16 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [65]. The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
E.2 In-domain prompt generation test
(a) Short Caption
(b) TIPO-Generated Caption
Figure 17: Comparison of generated images using original input (left) vs. TIPO-enhanced input (right)
Figure 17 illustrates the differences between short captions, truncated long captions, TIPO-generated captions, and TIPO-extended captions. The “short prompt” and “truncated long prompt” used in this experiment typically consist of 1-2 sentences, resulting in reasonably good quality outputs. However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
Appendix F Topic Distribution Visualization
Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [31], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words. LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [42] enhance its ability to extract meaningful insights from large textual datasets. We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus. This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [59].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [66]. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 7, 8, and 9. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 170.78 137.94 1078.38 204.71 125.90 144.59 1037.44 271.77
30 461.79 417.74 1758.40 736.74 327.48 195.13 1323.88 512.07
50 829.68 730.73 861.15 1036.33 400.90 373.74 823.51 959.18
100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
Table 7: Inertia for COYO-Dataset inference, higher is better
Size Promptdb GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 184.53 352.16 244.79 246.15 125.47 198.94 278.56 211.98
30 452.01 566.74 505.56 441.34 204.28 328.70 372.07 471.28
50 571.77 895.47 1227.30 990.17 438.89 313.48 737.65 788.41
100 1291.60 1742.36 1675.41 1550.32 631.61 628.78 1573.47 1855.90
Table 8: Inertia for GBC-Dataset inference, higher is better
Size PromptDB GPT4o-mini Promptist TIPO
20 60.82 734.52 210.60 139.29
30 275.76 1141.77 415.95 355.20
50 630.50 826.29 722.75 1002.36
100 2026.39 1879.08 802.93 1883.70
Table 9: Inertia for Scenery extend inference, higher is better
We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
Figure 18: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.
Figure 19: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), meaning it includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing tags that are both highly connected and influential within their network.