### v3

# TIPO: Text to Image with Text Presampling for Prompt Optimization

Abstract
TIPO (Text-to-Image Prompt Optimization) introduces an efficient approach for automatic prompt refinement in text-to-image (T2I) generation. Starting from simple user prompts, TIPO leverages a lightweight pre-trained model to expand these prompts into richer, detailed versions. Conceptually, TIPO samples refined prompts from a targeted sub-distribution within the broader semantic space, preserving the original intent while significantly improving visual quality, coherence, and detail. Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks.
We provide visual results, human preference report to investigate TIPO’s effectiveness. Experimental evaluations on benchmark datasets demonstrate substantial improvements in aesthetic quality, significant reduction of visual artifacts, and enhanced alignment with target distributions along with significant human preference proficiency. These results highlight the importance of targeted prompt engineering in text-to-image tasks and indicate broader opportunities for automated prompt refinement.
1 Introduction
The rapid proliferation of Text-to-Image (T2I) generative models have revolutionized artistic creation [48, 7, 20, 57, 53, 54, 60, 55, 50, 59, 15, 14, 35, 21, 8]. These models offer direct control over generative visual content via text prompts. To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout. However, the increasing complexity of prompts also forces users to iteratively refine their inputs, making high-quality T2I artwork accessible only to those with significant prompt engineering expertise.
Recently, efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality. A straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner [42]. Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models. This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts. A more effective approach is to train LLMs directly on prompt data collected from model users [5, 17]. While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs. Recent work [26] trains LLM with reinforcement learning, where aesthetic scores of generated images serve as rewards. However, the reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
Figure 1: Comparison of prompt optimization methods using LLM. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
Deviating from existing work, we argue that the optimal prompt lies in T2I models’ training set. Precisely, when a prompt aligns with the data distribution used for model training, the model can better interpret the prompt, resulting in improved text alignment and higher image quality. Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. At its core lies a lightweight language model trained on multiple meticulously designed pretext tasks that transform prompts into tags or sentences from coarse to fine. Leveraging the trained model, TIPO can progressively and seamlessly adapt to various user input types, generating accurate and robust prompts for various T2I models. In addition, our training data is derived from a curated text description dataset comprising over 30 million text-image pair samples and more than 20 billion tokens, ensuring comprehensive coverage of T2I training sets. Extensive experiments on test prompts from in- and out-of-domain demonstrate that TIPO consistently outperforms state-of-the-art methods in fidelity, aesthetics, and text alignment. Detailed analysis shows that TIPO has up to 62.80% win rate in user preference and 25% improvement in runtime efficiency over the second-best baseline. To summarize, our contributions are at least three-fold:
  1. 1.
    We introduce TIPO, a prompt optimization framework that leverages the large-scale text data distributions used in text-to-image (T2I) training.
  2. 2.
    We train a multi-task language model, which adaptively and progressively refines user inputs from either tag-based or natural language prompts into unified formats, enhancing compatibility across a broad spectrum of T2I models.
  3. 3.
    Extensive experiments demonstrate that TIPO achieves superior image quality, strong text alignment, higher human preference, and improved runtime efficiency across both in-domain and out-of-domain prompts, highlighting its great potential in real-world applications.
2 Related Work
Prompt optimization for T2I models typically leverages language models to refine user inputs. They can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that aim to improve prompt quality across a variety of T2I models.
2.1 Model-specific Strategies
T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference. These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model. For instance, Promptist [26] fine-tunes a pre-trained language model by using CLIP relevance scores as rewards. Similarly, PAE [44] extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning. However, these methods are computationally intensive, often struggling with a larger number of training prompts. Moreover, a model optimized for one specific T2I system may not generalize well to others. In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
2.2 Universal Strategies
To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models. For example, CogView3 [71] employ GLM-4 [24], and Lee et al. [33] employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement. However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models. Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs [5, 63, 17]. Such methods, however, are limited by the inconsistent expertise of users. Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
3 Preliminaries
Notations. We present the notations in our work:
  • $\mathcal{P}$: The space of all possible prompts.
  • $\mathcal{I}$: The space of all possible images.
  • $p$: User-provided prompt.
  • $p_{o}$: Optimized prompt.
  • $p_{ps}$: Sampled modified prompt derived from p.
  • $\mathcal{M}$: Metadata (e.g., style tags) aiding prompt transformation.
  • $\mathcal{T}$: Transformation function for pre-sampling.(e.g., Random augmentation, TIPO or LLM)
  • $T_{p},T_{ps}$: Sets of tags for original and pre-sampled prompts.
  • $S_{p},S_{ps}$: Ordered lists of natural language descriptions for original and pre-sampled prompts.
Text-to-Image Model. A text-to-image (T2I) model is a mapping function
$f(p)\;\;\to\;\;\mathcal{I}_{p}\subseteq\mathcal{I},$
which maps a user-provided text prompt $p$ to a subset of images $\mathcal{I}_{p}$ corresponding to the prompt.
Problem statement. Let $\mathcal{I}_{u}$ denote the user’s intended distribution over images. The task of Prompt optimization is to generate a prompt $p_{o}$ that minimizes a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and $\mathcal{I}_{u}$:
$p_{o}\;=\;\arg\min_{p\in\mathcal{P}}\;d\bigl{(}\mathcal{I}_{p},\,\mathcal{I}_{u}\bigr{)},$
where $d$ is a suitable metric (e.g., Frechet Inception Distance) comparing two image distributions.
[…]
4 Methodology
We aim to optimize user prompts to enhance image generation quality. Instead of end-to-end optimizations tailored to a single T2I model, our focus is on prompt rewriting that generalizes across a broad spectrum of models. Our core intuition is that an ideal prompt should align with the texts used in T2I model training. However, rapid advances in image captioning have rendered these texts increasingly diverse and complex. To address this, we propose to (1) design a clearly structured prompt schema compatible with most text descriptions, and (2) implement a pre-sampling algorithm that progressively refines arbitrary, coarse user input into organized, fine-grained prompts.
4.1 Text Set Preparation
Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions. Tag-based captions, such as those in the Danbooru2023 dataset [46, 68], use comma-separated, succinct terms to describe image content. In contrast, NL-based captions, typically generated by language models with visual capabilities [37, 1, 2, 34, 6, 16, 67, 18], may comprise multiple sentences. We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
While the original image captions are fine-grained and detailed, which can yield high-quality images when all elements are used, they often result in prompts that are excessively lengthy or overloaded with information. Such prompts diverge from typical user input and pose alignment challenges. To mitigate this, we construct a simpler subset $T_{s}\subset T$ by removing some tags and sentences from the original set [^1] For long sentences, off-the-shelf text splitters (e.g., FastText) are used to segment them into multiple sentences., as detailed in Section 4.2.
4.2 Formatted Prompt Construction
We aim to construct prompts in a unified format compatible with existing image captions. First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>. These metadata categories primarily include artist, copyright, aspect ratio, quality, and year (e.g., quality: masterpiece, artist: Picasso). This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
Next, we construct both tag-based and NL-based prompts using text sets $T$. Our design generates both simple (incomplete) and complete prompts for each image, and we train an auto-regressive language model to extend the simple prompts into complete versions. For tag-based prompts, since the tags are largely order-insensitive (i.e., the order has minimal impact on T2I outcomes), we propose a prefix-based dropout strategy. We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption. Then, we construct a simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags. The prompts are constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{o}=\text{concat}(T)$
Because $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand the simple tag-based prompt into its complete version.
For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information [25] and the order of sentences significantly influences the caption’s semantics. Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order. Let:
$S=[\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}]$
represent the ordered sequence of sentences in an image caption. We derive a simple subsequence $S_{s}$ by randomly selecting $m$ sentences from $S$ while ensuring that the first sentence is always included and that the original order is maintained. In other words,
$S_{s}=[\text{sentence}_{1},\text{sentence}_{i_{2}},\ldots,\text{sentence}_{i_{m}}],$
with $1<i_{2}<\ldots<i_{m}\leq n$ and $m<n$. The simple and complete NL-based prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{o}=\text{concat}(S_{s},S)$
This ensures that $p_{s}$ remains a prefix of $p_{o}$. Although some sentences may be repeated in $p_{o}$, selecting a smaller $m$ effectively mitigates this, and it does not empirically affect the generation quality.
4.3 Text Pre-sampling
We aim to align user input with the constructed high-quality training prompts $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling” - a process that first constrains the target distribution in text space before proceeding to image generation. Conceptually, while pre-sampling strategies generally add more text tokens to user input, the outcomes of different strategies may still vary (as illustrated in Figure 2).
To enrich the image content while mitigating undesired output, we perform pre-sampling based on a language model trained on $p_{o}$. A straightforward pre-sampling strategy is to do text completion, i.e., directly append output tokens to the end of user input. However, we posit that the user input is unlikely to be a part of a good prompt. The na’́ive text-completion strategy prioritizes the user’s original input, resulting in worse outcomes as evidenced by Section 5.6.
Figure 3: TIPO’s task flow within a single generation. Starting from any node, each arrow represents a sequential extension step, with other prompts used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
We propose the core technique of TIPO, a flexible pre-sampling mechanism that divides the prompt optimization process into three interconnected subtasks: enriching tag sequences, extending NL prompts, and refining NL prompts. As shown in Figure 3, TIPO uses both tag and NL prompts interchangeably as input sources to generate detailed outputs in various forms. For instance, TIPO can begin with a short NL prompt to produce a detailed tag sequence, denoted as short_to_tag. Training data for the language model is constructed by concatenating different types of $p_{o}$ with task-specific metadata embedded into the text.
The individual tasks are defined as follows:
  • tag_to_long: Use tags as metadata to generate a new NL prompt.
  • long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  • short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
  • short_to_long: Use a user-provided NL prompt as metadata to generate a refined, detailed prompt.
We also consider composite tasks that can be performed within a single forward pass. These tasks offer the language model a holistic view of the objective while reducing computational overhead:
  • short_to_tag_to_long: Use a user-provided NL prompt or tag sequence as metadata to generate a refined, detailed prompt.
  • short_to_long_to_tag: Use a user-provided NL prompt or generated NL prompt as metadata to extend a tag sequence.
  • tag_to_short_to_long: Use user-provided tags or NL prompts as metadata to generate a refined NL prompt.
We randomly select from the aforementioned tasks during training to enhance model generalization. By extensively training on these tasks, TIPO can seamlessly adapt to various input types, flexibly refining user input whether it consists of tags, short sentences, or long sentences. Figure 4 illustrates a scenario where both tag captions $T_{s}$ and short NL captions $S_{s}$ are available. In such cases, TIPO processes each input type separately to maintain clarity and coherence:
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\texttt{metadata}=\emptyset\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, \ldots}\right\}\end{array}$
The generation proceeds sequentially as follows:
  1. 1.
    short_to_tag: TIPO uses $T_{s}$ as the primary prompt to generate a detailed tag sequence $T_{d}$.
  2. 2.
    tag_to_long: $T_{d}$ is incorporated into the metadata, and TIPO produces a refined short NL prompt $S_{s}$ based on $T_{d}$.
  3. 3.
    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed output.
  4. 4.
    TIPO aggregates $T_{d}$, $S_{d}$, and any additional metadata to construct a context-rich prompt $p_{d}$.
This progressive process enables TIPO to build prompts that are both detailed and contextually aligned with the user’s input.
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
4.4 Implementation Details
TIPO can utilize various types of causal autoregressive language models by design. In practice, we leverage the LLaMA architecture [64, 65, 3] for simplicity, with 200M and 500M parameters.
  • TIPO-200M: Pretrained on the Danbooru2023 [46, 69] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
5 Experiments
5.1 Experimental Settings
Baselines. We compare with state-of-the-art prompt optimization methods as follows:
  • GPT [2] series are powerful large language models capable of zero-shot prompt optimization. We use GPT-4o-mini for lower cost and better efficiency.
  • MagicPrompt [17] trains GPT-2 on high-quality prompts collected from stable-diffusion model users.
  • Promptist [26] optimizes user input into model-preferred prompts via reinforcement learning.
Evaluation Metrics. We employ four latest metrics FDD [61], Aesthetic Score [19], AI Corrupt Score [45], and Vendi Score [22] to measure the quality of generated images. Specifically, FDD (Frechet Dino Distance) measures the image fidelity using the similarity between generated and dataset images via DinoV2 features [47], which better aligns with human perception than traditional FID [28]. Aesthetic Score is computed via Aesthetic Predictor V2.5 [19], quantifying visual appeal, composition quality, and artistic merit. AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts. Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings.
Evaluation Protocols. Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models. Following Promptist [26], we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
5.2 In-domain Tag-based Prompt Optimization
To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds. We randomly sample 32,768 tag-based prompts from Danbooru2023 [69, 46], shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta, an SDXL [50, 46] variant fine-tuned on Danbooru2023. This evaluation tests the in-domain capabilities, as Danbooru2023 is also used during TIPO training.
Metric Original GPT MagicPrompt Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Vendi Score $\uparrow$ 16.8135 8.6632 11.9014 14.3273 13.3070
Table 1: Performance on in-domain tag-based prompts, with best results bolded and the second-best underlined.
The results in Table 1 reveal two key insights. First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost). In contrast, GPT and TIPO benefit from large-scale training corpora ($>$30M samples), yielding higher-quality outputs. Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
5.3 In-domain NL-based Prompt Optimization
We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium [12] and GBC [29] as test prompts. In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences ($<$ 40 words) to simulate real-world applications. We use SDXL-1.0-base [50] as the T2I model, whose training text data largely overlap with TIPO.
Short Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Vendi Score $\uparrow$ 38.1715 34.7139 38.1553 34.1270 37.0649
Truncated Long Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Vendi Score $\uparrow$ 38.2533 34.8108 37.8412 33.5266 37.0900
Table 2: Performance on in-domain NL-based prompts.
(a) Scenery Tag Only
(b) Scenery Tag + TIPO
(c) Truncated Long Prompt
(d) Truncated Prompt + TIPO
Figure 5: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated ($<$ 40 words) long prompt, (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise. While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
(a) Oringial Prompt
(b) Text Completion
(c) TIPO
Figure 6: Qualitative comparison of prompt completion versus TIPO (ours). (a) shows images generated by original prompts. (b) depicts images generated by text completion using TIPO-500M. (c) shows images generated by TIPO.
5.4 Out-of-domain Performance
Some recent T2I models are trained on proprietary images and captions. As a representative example, SD-3.5-Large[20] is trained on private images captioned with CogView[71], which differ markedly from the texts used to train TIPO. To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets. We apply the remaining methods to these prompts and assess their performance.
GPT MagicPrompt Promptist TIPO(Ours)
Aesthetic Score $\uparrow$ 6.7125 6.4507 6.3924 6.0536
AI Corrupt Score $\uparrow$ 0.9482 0.8577 0.9053 0.9280
Vendi Score $\uparrow$ 8.9718 15.8721 16.4891 21.5706
Table 3: Performance on the out-of-domain T2I model
As shown in Table 3, SD-3.5-Large faithfully generates images that align well with GPT-produced prompts. Consequently, additional optimizations tend to reduce fidelity and introduce more artifacts. Nevertheless, GPT-generated prompts are accurate and lack diversity. TIPO optimization enriches the prompts with additional details that harmonize with the original themes, significantly enhancing the diversity of the generated images.
5.5 Human Preference Evaluation
Quantitative metrics may not fully align with human preference. Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers. As illustrated in Figure 7, TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors. In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts. For further results and statistics, please refer to Appendix F.
Figure 7: The win-tie-lose rate on overall preference on all data and out-of-domain situation where images are generated by SD-3.5-Medium.
Figure 8: ELO rating on overall preference on all data
5.6 Ablation Study
Original Completion TIPO
Aesthetic Score $\uparrow$ 5.9507 5.9631 5.9960
Corrupt Score $\uparrow$ 0.6799 0.8131 0.8971
Vendi Score $\uparrow$ 19.7208 18.2896 13.9864
Table 4: Ablation results
We attribute TIPO’s success primarily to its multi-task pre-sampling design detailed in Section 4.3. However, a natural question arises: can a simple text completion approach achieve similar or better results? To validate our design, we compare TIPO’s pre-sampling strategy with a direct text completion using the same language model, TIPO-500M. As shown in Table 4, TIPO significantly improves aesthetic scores and reduces artifacts compared to text completion. For visual examples, please refer to Figure 6. For more ablation details and visualization, please refer to Appendix 5.6.
6 Conclusion
We introduced TIPO, a lightweight prompt pre-sampling framework designed for efficient real-world Text-to-Image (T2I) applications. By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead.
TIPO consistently outperforms existing prompt optimization methods, achieving superior performance across multiple evaluation metrics. Its ability to preserve user intent while enhancing prompt diversity and image quality makes it well-suited for both in-domain and out-of-domain scenarios, demonstrating robustness across various prompt formats and model architectures.
Extensive user studies also confirmed TIPO’s strong alignment with human preferences, further highlighting its potential for practical applications. To encourage wider adoption and facilitate reproducibility, we release our trained models and source code. We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.
Appendix
[…]
Appendix A Dataset/Resource
A.1 Danbooru2023
The Danbooru2023 dataset [69, 68, 46] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing various visual content with granular annotations. These annotations allow for creating nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.
Key Characteristics:
  • Rich Annotations: Detailed tags differentiate subtle variations, crucial for specific image generation.
  • Large Volume: Extensive dataset size ensures diverse training examples.
  • Tag-Based Prompting: Refined prompts from detailed tags enhance image generation accuracy.
A.2 GBC10M
The GBC10M dataset [29] is a large-scale collection of 10 million images sourced from CC12M [13], annotated using the Graph-Based Captioning (GBC) approach. Each image is represented by a graph where nodes correspond to object regions, compositions, and relations, and edges define their hierarchical relationships. Annotations are generated automatically through a pipeline leveraging pretrained multimodal large language models (MLLM) and object detection tools. The GBC structure enhances traditional image captions by providing detailed descriptions and structural information. Data is provided in JSON lines format, including image URLs, bounding boxes, and captions.
In TIPO, only the root node captions from GBC10M are utilized for concise yet descriptive prompts.
A.3 Coyo HD 11M
The Coyo HD 11M dataset [12] consists of 11.4 million high-resolution, high-concept-density images paired with 22.8 million synthetic captions generated from the Coyo-700M dataset. Images maintain a minimum of 512 pixels on the shortest edge to ensure high visual quality. Captions, generated with the LLaVA-Next-8B model [38] based on LLaMA 3 [3], undergo post-processing for conciseness and clarity.
TIPO uses short and long captions, booru tags, and open image tags from this dataset.
A.4 Stable Diffusion XL
Stable Diffusion XL (SDXL) [50] improves upon earlier models [56] with a more considerable UNet backbone and dual text encoders (CLIP ViT-L [51] and OpenCLIP ViT-bigG [31]), enhancing text conditioning. Supporting resolutions up to 1024×1024, SDXL accepts natural language prompts and tags, suitable for diverse image generation.
A.5 Illustrious
Illustrious is a series of fine-tuned Stable Diffusion XL models primarily trained on the Danbooru2023 dataset. In this study, we specifically employ the v3.5 version variant with v-parameterization [58], which is notable for its extensive incorporation of natural language prompts. The inclusion of both tag-based and natural language formats allows Illustrious to leverage a broad range of semantic knowledge for image generation.
Within TIPO, we perform an ablation study to analyze the effectiveness of different prompting strategies—namely extended tags versus natural language prompts—to identify which approach contributes most significantly to enhanced image generation performance.
A.6 Stable Diffusion 3.5
Stable Diffusion 3.5 (SD-3.5) incorporates the MMDiT architecture [20] and the Rectified Flow formulation [39, 4, 36] for improved text-to-image generation. Utilizing triple text encoders (CLIP/ViT-L, OpenCLIP/ViT-G, T5-XXL [52]), SD-3.5 supports resolutions up to 1024×1024 and uses a 50/50 mix of original and CogVLM-generated captions. Figures confirm the capability to process both natural language prompts and tags.
This study employs SD-3.5-Large [^6] https://huggingface.co/stabilityai/stable-diffusion-3.5-large (8B parameters) with FP8 inference on RTX 3090 or RTX 4090 GPUs.
Appendix B Implementation Details
In this appendix, we provide all the necessary details including our dataset construction process, model configurations, inference pipeline and the model’s properties not mentioned in Section 4.2 and 4.3.
B.1 Training Data Construction
This section details our methodology for constructing and preprocessing training data to ensure robust model performance across various input scenarios.
Length Control
To systematically control output prompt length, we implement a structured length categorization system using unique length tags. These tags enforce specific constraints on tag counts and natural language sentence lengths. For instance, the <long> tag specifies that the corresponding prompt must contain between 36 and 52 tags (inclusive), accompanied by 4 to 8 sentences of natural language description. We define four distinct length categories, each with strict bounds for tag count and sentence length.
Type Very Short Short Long Very Long
Tags (count) 18 36 48 72
NL (sentences) 2 4 8 18
Table 5: Maximum length specifications for each category and caption type. For each category, the actual count/length must not exceed these values.
Random Augmentation
To enhance input diversity and better simulate real-world usage patterns, we implement several data augmentation strategies:
  • Metadata Tags: For tags representing image metadata (e.g., artist, character, aspect ratio), we employ two randomization techniques:
    • Random removal of metadata tags
    • Random repositioning of metadata tags to the end of the prompt, after all content-related descriptions
    This approach encourages the model to handle varying metadata positions and availability, while maintaining the ability to infer metadata relationships from content descriptions.
  • Content Tags: For tags describing image content (e.g., objects, actions, attributes), we implement:
    • Random shuffling of tag order within the content section
    • Length-based truncation to meet target length constraints while preserving key content information
  • Natural Language: For natural language descriptions exceeding length limitations, we employ selective sentence removal, targeting middle sentences to preserve context-setting opening sentences and concluding details. This maintains coherent narrative flow while meeting target length requirements.
These augmentation strategies create a more diverse training dataset that better reflects real-world prompt variations, improving the model’s robustness and adaptability to different input styles and formats.
B.2 Training Settings and Model Configurations
TIPO-100M TIPO-200M stage1 TIPO-200M stage2 TIPO-500M
Architecture LLaMA
Type Pretrain Pretrain Finetune Pretrain
Vocab Size 32013
Hidden Dim 640 768 - 1280
Attention Heads 10 12 - 20
MLP Dim 2240 2304 - 3840
Hidden Layers 10 20 - 20
Model Parameters 100M 203M - 508M
Max Learning Rate 5e-4 2e-4 5e-5 2e-4
Optimizer AdamW
LR scheduler Cosine Annealing LR
betas 0.9, 0.98
weight decay 0.01
Dataset Coyo, GBC, Dan GBC, Dan Coyo, GBC, Dan Coyo, GBC, Dan
Epoch 1 5 3 5
max context length 512 512 1024 1024
global batch size 1024 2048 2048 3584
Token Seen 6.0240B 22.625B 18.339B 31.274B
Hardware 4 × RTX3090 4 × RTX3090 4 × RTX3090 8 × H100
Training Time (wall) 22.5 hour 150 hour 270 hour 100 hour
Table 6: Training settings for TIPO models. The datasets include CoyoHD11M (Coyo), GBC10M (GBC), and Danbooru2023 (Dan). Stage 2 additionally incorporates Pixtral [1] to generate NL captions from Danboodu2023 dataset.
Tokenizer and Task Tokens
TIPO employs a vocabulary derived from LLaMA2 [65] consisting of 32,000 tokens, with additional tokens (13 tokens) specifically designated for task and length control or placeholders. This extended vocabulary includes task identifiers and length modifiers to ensure flexibility across different prompt types:
  • Placeholder Token (1 token):
        <|empty|>
  • Task Tokens (8 tokens):
        <|gen_meta|>, <|tag_to_long|>, <|short_to_tag|>,
        <|long_to_tag|>, <|short_to_long|>, <|short_to_tag_to_long|>,
        <|short_to_long_to_tag|>, <|tag_to_short_to_long|>
  • Length Tokens (4 tokens):
        <|very_short|>, <|short|>, <|long|>, <|very_long|>
Optimizer and Learning Schedule
Training is performed using the AdamW optimizer [40], with a cosine annealing learning rate scheduler [41]. The optimizer parameters include $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of 0.01. Maximum learning rates are adjusted per model size, as outlined in Table 6.
Training Configurations
TIPO models are trained in multiple stages. Table 6 summarizes the configurations for pretraining and fine-tuning TIPO-100M, TIPO-200M, and TIPO-500M. Both pretraining and fine-tuning was conducted on datasets like Danbooru2023 [46], GBC10M [29], and CoyoHD11M [12].
Augmented Task Representation
Each dataset entry undergoes random task assignment and splitting to simulate a wide range of input-output mappings, effectively increasing the dataset size. For example, a single entry may contribute to tasks like short_to_tag or tag_to_long, with length modifiers dynamically controlling the output verbosity. This approach ensures the model can handle diverse tasks while maintaining robust generalization.
Hardware and Time Requirements
Training was conducted on NVIDIA RTX3090 GPUs for smaller models and H100 GPUs for TIPO-500M. Total wall-clock training times ranged from 22.5 hours for TIPO-100M to 270 hours for fine-tuning TIPO-200M.
Token Seen and Effective Training
Non-padding tokens are used to measure the effective token count during training, ensuring efficiency given the short and variable data lengths. Table 6 details the total tokens seen per model and training stage, illustrating the comprehensive exposure to diverse data entries.
[…]
Figure 9: TIPO inference workflow, with solid arrows denoting the primary generation steps and dashed arrows indicating alternative generation paths within the same cycle. ¡TOKEN¿— represents special tokens, with all tokens detailed in Section B.2.
B.3 Inference Pipeline
The TIPO inference pipeline is designed to handle various input types and scenarios, combining different tasks to refine or expand both tag-based and natural language prompts. Figure 9 illustrates this comprehensive workflow.
Our framework processes tags and natural language inputs separately, allowing for specialized handling of each input type. This flexible pipeline allows TIPO to adapt to various input scenarios, whether the user provides tags, natural language descriptions, or both. By leveraging different task combinations, TIPO ensures that tag-based and natural language prompts are optimized, resulting in more detailed and effective input for text-to-image models.
B.4 Model Speed Comparison
We conducted comprehensive speed tests of our 100M, 200M and 500M parameter models using the inference pipeline described in Section 5.2. Each prompt requires two sequential generation steps. Our primary metric is average tokens generated per second, which reflects real-world task performance rather than theoretical maximum throughput.
The evaluation was performed using llama.cpp [23], an efficient C++ implementation that provides optimized support for various hardware accelerators including CUDA, HIP, and Apple Metal.
TIPO-100M TIPO-200M TIPO-500M
Hardware Platform tok/sec gen time tok/sec gen time tok/sec gen time
M1 Max (32 GPU cores) 339.4 0.66 190.0 1.23 119.4 2.02
RTX 3090 558.5 0.42 341.4 0.69 289.8 0.81
RTX 4090 742.9 0.29 454.5 0.51 359.7 0.63
Table 7: Model performance comparison across different hardware platforms. Tokens per second (tok/sec) represents the average generation speed, while generation time (gen time) shows the average time in seconds required for a complete two-step prompt optimization process.
Appendix C Evaluation Statistics
In this appendix, we provide more statistics for the result obtained in Section 5.
C.1 In-domain test regarding scenery tag
(a) The box plot for the Aesthetic Score result of scenery tag test.
(b) The box plot for the AI Corrupt Score result of scenery tag test.
(c) The KDE plot for the Aesthetic Score result of scenery tag test.
(d) The KDE plot for the AI Corrupt Score result of scenery tag test.
Figure 10: The distribution of aesthetic and AI corrupt score for scenery tag test.
The box plot and Kernel Density Estimation (KDE) plot displayed in Figure 10 illustrate the aesthetic scores and AI corruption scores from the scenery tag test described in Section 5.2. The analysis shows that TIPO significantly outperforms all other methods, demonstrating a considerable margin of improvement.
C.2 In-domain prompt generation test
(a) The box plot for the Aesthetic Score result of short prompt input.
(b) The box plot for the AI Corrupt score result of short prompt input.
(c) The KDE plot for the Aesthetic Score result of short prompt input.
(d) The KDE plot for the AI Corrupt Score result of short prompt input.
Figure 11: The distribution of aesthetic and AI corrupt score for short prompt input.
(a) The box plot for the Aesthetic Score result of truncated long prompt input.
(b) The box plot for the AI Corrupt score result of truncated long prompt input.
(c) The KDE plot for the Aesthetic Score result of truncated long prompt input.
(d) The KDE plot for the AI Corrupt Score result of truncated long prompt input.
Figure 12: The distribution of aesthetic and AI corrupt score for truncated long prompt input.
Figures 11 and 12 display the box plots and KDE plots of aesthetic scores and AI corruption scores obtained from the In-domain prompt generation test detailed in Section E.2. While the box plots reveal subtle differences in performance between various methods, the AI corruption scores provide valuable insights. Specifically, these scores indicate that implementations supported by TIPO produce more stable output images than other methods.
C.3 Out-of-domain evaluation
(a) The box plot for the Aesthetic Score result of out-of-focus test.
(b) The box plot for the AI Corrupt Score result of out-of-focus test.
(c) The KDE plot for the Aesthetic Score result of out-of-focus test.
(d) The KDE plot for the AI Corrupt Score result of out-of-focus test.
Figure 13: The distribution of aesthetic and AI corrupt score for out-of-focus test.
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by MagicPrompt
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 14: The similarity matrix for the 100 best aesthetic results generated in the SD3.5-Large experiments. Off-diagonal elements of the matrix indicate the similarity between different images. A lower value for an off-diagonal element indicates greater diversity among the generated images.
(a) Original Caption
(b) Prompt by GPT4o-mini
(c) Prompt by MagicPrompt
(d) Prompt by Promtist
(e) Prompt by TIPO
Figure 15: The similarity matrix between 100 images of worst aesthetic generated results of SD3.5-Large experiments.
Figures 14 and 15 present similarity matrices for different prompt generation methods and their corresponding aesthetic outputs on SD3.5-Large  [21]. A matrix with predominantly lower similarity values (brighter appearance) indicates high diversity among generated images, while higher values (darker appearance) suggest consistent but less diverse outputs. Please refer to Table 3 in Section 5.
C.4 Ablation Test
(a) The box plot for the Aesthetic Score result of NL ablation test.
(b) The box plot for the AI Corrupt Score result of NL ablation test.
(c) The KDE plot for the Aesthetic Score result of NL ablation test.
(d) The KDE plot for the AI Corrupt Score result of NL ablation test.
Figure 16: The distribution of aesthetic and AI corrupt score for NL ablation test.
(a) The box plot for the Aesthetic Score result of tags ablation test.
(b) The box plot for the AI Corrupt Score result of tags ablation test.
(c) The KDE plot for the Aesthetic Score result of tags ablation test.
(d) The KDE plot for the AI Corrupt Score result of tags ablation test.
Figure 17: The distribution of aesthetic and AI corrupt score for tags ablation test.
Figures 17 present the tag ablation test in the TIPO effect on the aesthetic score and AI Corrupt Score among the original tag, tag-extend and the tags TIPO. The box plot reveals that the tag TIPO is better than the original tag and the tag extend is the best. In detail, KDE plot reveals that the tag TIPO has a similar performance compared with the tag extend. Both of them are better than the original tag, which indicates that the tag TIPO aspect helps control corruption and promotes the aesthetic score.
Appendix D TIPO example
In this section, we provide some text example of TIPO’s input and output.
TIPO Format Template User Input: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, masterpiece, absurdres, newest, safe A girl sits in a cozy cafe, cradling a cup of coffee in her hand Formatted TIPO Input for Expand Tags: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|short_to_tag|> short: A girl sits in a cozy cafe, cradling a cup of coffee in her hand tag: 1girl, solo Formatted TIPO Output after Expand Tags and Expand Natural Prompt: meta: absurdres rating: safe artist: ciloranko, maccha (mochancc), ningen mame, ask (askzy) quality: masterpiece, newest aspect_ratio: 1.0 target: <|short|> <|tag_to_long|> tag: 1girl, solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding long: A girl sits in a cozy cafe, cradling a cup of coffee in her hand. The cafe has large windows with green plants on the walls and a wooden table in front of her. Behind her is a staircase leading to another room. The overall atmosphere of the image is serene and inviting. Formatted Output for Text-to-Image: 1girl, ciloranko, maccha (mochancc), ningen mame, ask (askzy), solo, sitting, closed mouth, jewelry, long hair, looking at viewer, crossed legs, plant, table, couch, bracelet, cup, smile, teacup, indoors, blue eyes, blonde hair, holding, A girl sits in a cozy cafe, cradling a cup of coffee in her hand. Behind her is a staircase leading to another room. The cafe has large windows with green plants on the walls and a wooden table in front of her. The overall atmosphere of the image is serene and inviting. masterpiece, newest, absurdres, safe Refer to caption User Input Refer to caption TIPO Output
Figure 18: An example of formatted content used for training and inference in TIPO.
TIPO Format template User Input:
scenery, no humans, masterpiece, absurdres, newest, safe
Formatted TIPO Input For Expand Tags:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> tag: scenery, no humans
Formatted TIPO Output after Expand Tags and tag_to_long task:
meta: absurdres rating: safe quality: masterpiece, newest aspect_ratio: 1.0 target: <|long|> <|tag_to_long|> tag: scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar (cafe), lamp, road long: A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. on top of the traffic light pole, there seems to be an air conditioning unit. the street is lined with trees and bushes, and there is graffiti on the ground.
Formatted Output For Text-to-Image
scenery, no humans, storefront, motor vehicle, road sign, power lines, plant, railing, flower pot, vanishing point, outdoors, sign, potted plant, sidewalk, awning, tree, bicycle, window, railroad crossing, bush, building, utility pole, lamppost, shop, truck, traffic light, fence, chinese text, stairs, door, bicycle basket, town, day, streetcar \(cafe\), lamp, road, A small town with a variety of buildings and houses. the sky is blue and there are trees in the background. on the left side of the image, there is an orange building with a sign that reads "chinese restaurant". on the right side, there are several other buildings with different types of shops and restaurants. in front of the buildings, there appears to be a street with cars parked along the road. in the center of the illustration, we can see a train crossing signal with two red lights and a blue sky above it. there is also a yellow building with white walls and a green roof. masterpiece, newest, absurdres, safe
Refer to caption User InputRefer to caption TIPO Output
Figure 19: An example formatted content we used for training and inference in TIPO.
Appendix E Image Examples
In this section, we present sample images from the experiments described in Section 5 to visually demonstrate the improvements achieved by TIPO.
E.1 In-domain test regard to scenery tag
Figure 20: Comparison of generated images using simple input (left) vs. TIPO-enhanced input (right) for the scenery tag
Figure 20 demonstrates the difference in output diversity between simple input and TIPO-enhanced input for the scenery tag. As observed, TIPO significantly expands the range of generated sceneries, better reflecting the variety present in the Danbooru2023 dataset [69]. The left column shows results from simple input (scenery tag only), while the right column illustrates the enhanced diversity achieved with TIPO-enhanced input.
E.2 In-domain prompt generation test
(a) Short Caption
(b) TIPO-Generated Caption
Figure 21: Comparison of generated images using original input (left) vs. TIPO-enhanced input (right)
Figure 21 illustrates the differences between short captions, truncated long captions, TIPO-generated captions, and TIPO-extended captions. The “short prompt” and “truncated long prompt” used in this experiment typically consist of 1-2 sentences, resulting in reasonably good quality outputs. However, the use of TIPO to refine or extend these prompts still yields noticeable improvements in aesthetics and overall quality.
Appendix F Human Preference
(a) ELO rating on Illustrious
(b) Win rate matrix on Illustrious
(c) ELO rating on SD3.5-medium
(d) Win rate matrix on SD3.5-medium
Figure 22: ELO ratings and win rate matrices across different experimental settings comparing five prompting methods (TIPO, Promptist, Promptextend, MagicPrompt, and Original) on three evaluation dimensions
(a) Full Win-Tie-Lose plot
(b) Illustrious Win-Tie-Lose plot
(c) SD3.5-medium Win-Tie-Lose
Figure 23: Win-Tie-Lose comparison across different experimental settings showing the relative performance of five prompting methods on prompt adherence, image quality, and aesthetic appeal metrics
(a) The UI of survey system before submitting the choices.
(b) The UI of survey system after submitting the choices.
Figure 24: Survey interface for human evaluation of image pairs, showing the evaluation process before submission (a) where users compare two images based on four metrics, and after submission (b) where the generated prompts for each image are revealed
Comparison Win Ratio (A:B) Proportion A $p$-value Significant?
Original vs. PromptExtend 45:66 0.405 0.0572 Marginally
Original vs. Tipo 28:80 0.259 $<0.0001$ Yes***
Original vs. Promptist 35:64 0.354 0.0046 Yes**
PromptExtend vs. Promptist 41:80 0.339 0.0005 Yes***
Promptist vs. Tipo 30:101 0.229 $<0.0001$ Yes***
PromptExtend vs. Tipo 30:90 0.250 $<0.0001$ Yes***
MagicPrompt vs. Promptist 14:34 0.292 0.0055 Yes**
MagicPrompt vs. Tipo 11:48 0.186 $<0.0001$ Yes***
MagicPrompt vs. Original 15:28 0.349 0.0660 No
MagicPrompt vs. PromptExtend 19:35 0.352 0.0402 Yes*
Table 8: Pairwise win rates and statistical significance (Overall Dimension). Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$
We conducted a series of A/B tests to compare five prompt transformations, (TIPO, Promptist, Promptext, MagicPrompt, and Original(unmodified)), for two models, Illustrious, SD3.5-medium, which is known for both core word/natural language understanding. In total, we collected responses for $\sim$1,500 pairwise comparisons, from more than 20 anonymous evaluators. Each evaluation asked participants to compare two generated images—labeled A and B—and select which they preferred (or a tie) according to specific criteria (e.g., prompt adherence, image quality, or aesthetic appeal).
F.1 User Interface for Human Preference Evaluation
We developed a specialized survey interface to facilitate efficient and unbiased human evaluation of generated images. As illustrated in Figure 24, the interface presents evaluators with an original prompt and two corresponding images (labeled A and B) generated using different prompting methods. Before submission, users can see the original prompt in the center panel while the processed prompts used to generate each image remain hidden to prevent bias.
The evaluation framework requires participants to compare the image pairs across four distinct metrics: prompt adherence (how well the image follows the original prompt), image quality (detail and correctness), aesthetic appeal (color, composition, and style), and overall personal preference. For each metric, users can select one of three options: “A is better,” “A and B are equal,” or “B is better.”
When evaluators encounter image pairs that appear to be from different prompts or settings, they are instructed to click “Refresh” to obtain a new comparison. After submitting their evaluations, the interface reveals the transformed prompts used to generate each image, providing transparency about how the original prompt was modified by each method.
F.2 Extended Human Evaluation.
Participants assessed each image’s performance on prompt adherence, image quality, and aesthetic appeal, with visually shown unmodified and image pairs. TIPO exhibited superior outcomes in all comparison settings. Notably, it attained a 64.4% peak win rate (against MagicPrompt) under the Full scenario and 57.5% (also against MagicPrompt) under SD35-medium, emphasizing TIPO’s proficiency in generating images that closely follow prompt specifications while maintaining visual coherence.
F.3 ELO Ratings.
We computed theoretical ELO ratings from the aggregated pairwise comparisons to quantify overall performance differences among the five methods. The rating update rules were based on each pair’s binary outcome (win or lose), ignoring tie cases. The result is depicted in Figure 22, TIPO has secured the highest ELO rating over other models.
F.4 Human Preference ELO Method
We computed theoretical ELO ratings from human-judged pairwise preference data to quantitatively evaluate the relative performance of each prompting method. The ELO rating system, initially designed for ranking chess players, aggregates binary outcomes into numerical ratings representing comparative performance.
Pairwise Outcomes.
Human evaluators assessed comparisons between methods, resulting in one of three outcomes:
  • Method $i$ wins: assigned a score of $1$ for method $i$, and 00 for method $j$.
  • Method $j$ wins: assigned score $1$ for method $j$, and 00 for method $i$.
  • Tie: assigned score $0.5$ to both methods.
Conversion to ELO Differences.
Win and tie rates were converted to ELO rating differences using:
Adjusted Win Rate=Win Rate+Tie $\text{Adjusted Win Rate}=\text{Win Rate}+\frac{\text{Tie Rate}}{2}$
ELO $\text{ELO Difference}=400\times\log_{10}\left(\frac{\text{Adjusted Win Rate}}{1-\text{Adjusted Win Rate}}\right)$
To ensure numerical stability, extreme adjusted win rates were constrained as follows:
ELO $\text{ELO Difference}=\begin{cases}-800,&\text{Adjusted Win Rate}\leq 0.001\\ +800,&\text{Adjusted Win Rate}\geq 0.999\\ \end{cases}$
Calculating Method ELO Ratings.
Final ELO ratings were determined by averaging each method’s pairwise ELO differences and centering these averages around a baseline rating (e.g., $1000$):
$\text{ELO}_{\text{method}_{i}}=\text{Base Rating}+\left(\text{Average ELO Difference for method }i-\text{Overall Mean ELO Difference}\right)$
Interpretation of ELO Scores.
Methods with higher ELO scores consistently outperform lower-scored methods. A rating difference of 400 points corresponds to a 90% expected win probability for the superior method.
F.5 Statistical Significance
As summarized in Table 8, we conducted two-sided binomial and McNemar’s tests (p <0.05) to assess the statistical significance of observed differences. The result confirms that TIPO’s advantages are unlikely to be explained by random variation, which also supports a consistent performance hierarchy: TIPO ranks highest, followed by Promptist, PromptExtend, Original, and MagicPrompt. Collectively, these findings illustrate TIPO’s robust, model-agnostic effectiveness and underscore the model-sensitivity of alternative methods, particularly Promptext and MagicPrompt.
F.6 Survey Response Examples
In this section we provided some responses of our human preference survey as reference.
Figure 25: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
Figure 26: Some survey responses on illustrious-3.5-vpred generated image with different prompt optimization method
Figure 27: Some survey responses on SD3.5-medium generated image with different prompt optimization method
Figure 28: Some survey responses on SD3.5-medium generated image with different prompt optimization method
F.7 Conclusion
The extended evaluations presented here reinforce TIPO’s standing as a reliable and effective prompt-optimization strategy. Its consistent performance gains under diverse model conditions highlight its potential for broad applicability, with its strong alignment with user-specified prompts, high image quality, and favorable aesthetic outcomes.
Appendix G Ablation Study on TIPO
In this section, we investigate the effect of incorporating TIPO (Tags + Inferred Prompt Objects) across various generation settings. Our primary goal is to validate whether additional structured information (e.g., core tags and minimal spatial/contextual cues) can improve image quality, reduce artifacts.
G.1 Experimental Setup
Prompt Variants.
To systematically analyze TIPO’s contribution, we consider four types of input prompts improvement task:
  1. 1.
    Tag $\to$ More core words: Given an initial set of core words, generate more refined or expanded core words.
  2. 2.
    NL $\to$ More NL: Given a short natural language (NL) description, elaborate into a richer NL prompt.
  3. 3.
    Tag $\to$ (More core words + NL): Combine expanded tags with a corresponding NL description derived from them.
  4. 4.
    NL $\to$ (More NL + core words): Use the NL prompt to add relevant tags, forming a mixed prompt of NL plus core words.
In each case, we compare the baseline prompts (without TIPO cues) against prompts incorporating TIPO’s structured, tag-based critical information and minimal spatial hints.
Data Preparation.
We start by randomly sampling core words from a word table to represent a diverse range of topics (e.g., objects, environments, descriptors). Additionally, for each word set, we generate a corresponding short NL sentence using a compact language model (GPT4o-mini). Overall, the six prompt variants are tested on 4,000 images, ensuring a balanced comparison.
Inference Procedure.
Prompts are fed into our image-generation pipeline under identical model settings (classifier free guidance, sampler, steps, etc.), using the v-parameterized variant of Illustrious v3.5[49]. We focus on how TIPO modifications alter the generation outcomes and whether they introduce additional computational overhead.
G.2 Evaluation Metrics
(a) Corrupt Score Distributions
(b) Aesthetic Score Distributions
Figure 29: Side-by-side comparison of Corrupt (left) and Aesthetic (right) score distributions across prompt types.
Aesthetic Score.
We employ an off-the-shelf aesthetic predictor to estimate image quality. In the following paragraph, we discuss the model’s bias.
AI Corruption Score.
Using an automated ’ AI corruption ’ detection model, we measure generation artifacts, such as distorted objects and unnatural shapes. Higher scores imply cleaner, more coherent outputs.
G.3 Results & Discussion
Impact on Aesthetics.
Figure 29(b) shows that TIPO-enhanced prompts generally achieve higher aesthetic scores than their non-TIPO counterparts, albeit with some variance. Notably, we observe a correlation between wider color ranges and higher aesthetic scores discussed in Figure 30, suggesting a bias toward more colorful or varied compositions.
Improvements in Corruption Score.
As shown in Figure 29(a), TIPO-based prompts yield significantly lower corruption scores, indicating fewer artifacts. We hypothesize that the additional spatial and contextual details encoded via TIPO help the model place objects more consistently.
(a) Saturation vs. Aesthetic Score
(b) Binned Saturation vs. Aesthetic
(c) Saturation vs. Corrupt Score
Figure 30: Scatter plots (left and right) and binned analysis (middle) showing the relationship between saturation and image metrics. We find a moderate positive correlation between saturation and aesthetic score (Pearson $r=0.2821$), particularly at lower saturation ranges, based on 24k samples. However, saturation shows no notable correlation with corrupt score (Pearson $r=0.0125$).
G.4 Speed Test and Overhead Analysis
A key concern for production pipelines is whether TIPO generation imposes a substantial time overhead. We benchmarked prompt-generation inference on four smaller models, excluding any large proprietary LLMs. As illustrated in Table 9, the additional TIPO-related computation remains well below the image-generation time. Hence, even in a synchronous pipeline, TIPO prompt expansion does not constitute a bottleneck.
Table 9: Speed Test Results for TIPO and Other Prompt Methods
Method Model/Config # Runs Avg. Time (s) Std. Dev. (s)
TIPO LLaMA-500M 500 1.4207 1.0730
LLaMA-200M 200 1.0306 0.8982
LLaMA-100M 200 1.0078 0.9394
PROMPTIST GPT2-125M 1000 1.4593 0.2857
PROMPTEXTEND GPT2-125M 1000 1.3849 0.2151
MAGICPROMPT GPT2-125M 1000 1.1398 0.4043
Memory Footprint.
We also confirm that TIPO’s overhead in terms of VRAM usage is minimal. The practical adoption of TIPO in pipelines has shown no critical memory concerns, which aligns with our measurements.
G.5 Conclusion of Ablation
Our experiments suggest that TIPO (1) consistently lowers AI corruption artifacts, (2) can boost aesthetic scores, and (3) remains computationally inexpensive. The improvements in metrics support the viability of TIPO prompts for real-world image-generation tasks. In short, a concise natural language prompt with core tag-based critical information appears to be an effective, suggested form for most use cases.
Appendix H Topic Distribution Visualization
Latent Dirichlet Allocation (LDA) [9] is a generative probabilistic model for topic modeling [32], which assumes that each document is a mixture of topics, with each topic represented by a distribution over words. LDA uncovers hidden thematic structures by analyzing word co-occurrence patterns, while methods like TF-IDF and TextRank [43] enhance its ability to extract meaningful insights from large textual datasets. We implemented a multi-stage topic modeling and clustering methodology using LDA to extract varying numbers of topics (20, 30, 50, and 100) from the corpus. This approach focuses on identifying significant representative words while filtering out stop words and irrelevant terms to ensure meaningful topic classification.
We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [62].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [70]. We grouped the initially extracted topics into five major clusters using k-means clustering. We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 10, 11, and 12. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Size MagicPrompt GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 170.78 137.94 1078.38 204.71 125.90 144.59 1037.44 271.77
30 461.79 417.74 1758.40 736.74 327.48 195.13 1323.88 512.07
50 829.68 730.73 861.15 1036.33 400.90 373.74 823.51 959.18
100 1656.74 1245.60 1987.63 1628.32 877.36 657.14 1622.79 1777.61
Table 10: Inertia for COYO-Dataset inference, higher is better
Size MagicPrompt GPT4o-mini Promptist TIPO
Run 1 Run 2 Run 1 Run 2 Run 1 Run 2 Run 1 Run 2
20 184.53 352.16 244.79 246.15 125.47 198.94 278.56 211.98
30 452.01 566.74 505.56 441.34 204.28 328.70 372.07 471.28
50 571.77 895.47 1227.30 990.17 438.89 313.48 737.65 788.41
100 1291.60 1742.36 1675.41 1550.32 631.61 628.78 1573.47 1855.90
Table 11: Inertia for GBC-Dataset inference, higher is better
Size MagicPrompt GPT4o-mini Promptist TIPO
20 60.82 734.52 210.60 139.29
30 275.76 1141.77 415.95 355.20
50 630.50 826.29 722.75 1002.36
100 2026.39 1879.08 802.93 1883.70
Table 12: Inertia for Scenery extend inference, higher is better
We attach a simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
Figure 31: Topic visualization for scenery prompt generation. A wider spread indicates a greater diversity of generated topics.
Figure 32: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), which includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing highly connected and influential tags within their network.
Appendix I Discussion and Future Work
Despite the promising performance demonstrated by TIPO, several areas merit further investigation:
Personalization and User Adaptation.
TIPO currently does not directly incorporate personalization methods (e.g. LoRA [30]). Incorporating such modules could further align generated outputs to individual user preferences, enabling more tailored creative assistance. Investigating joint training strategies that integrate personalized fine-tuning represents a valuable future direction.
Generalization to Uncommon Compositions.
While TIPO performs effectively on typical and broadly represented concepts, it may still struggle with highly unconventional combinations (e.g., “avocado chair”). Future research should systematically evaluate TIPO’s performance on rare or unusual prompts (but considered creative prompts), potentially exploring auxiliary training or adaptation methods to ensure broader compositional robustness.
Capacity and Backbone Scalability.
Given TIPO model’s architecture [66, 65], its capacity for handling highly complex or lengthy prompts could be inherently limited. Future work could investigate more extensive or specialized architectures enhance compositional and semantic capacity, ideally without compromising real-time inference efficiency, are crucial for practical deployment.